### v2

Abstract
Local superlinear convergence of the semismooth Newton method usually necessitates assumptions on the uniform invertibility of the utilized, generalized Jacobian matrices, such as, e.g., BD- or CD-regularity. For certain composite-type problems and nonlinear programs (for which explicit representations of the generalized Jacobians of the associated stationarity equations are available), such regularity assumptions are closely connected to strong second-order sufficient conditions. However, general characterizations are not well understood. In this paper, we investigate a strong second-order sufficient condition ($\mathrm{SSOSC}$) for composite problems whose nonsmooth part has a generalized conic-quadratic second subderivative. We discuss the relationship between the $\mathrm{SSOSC}$ and other second order-type conditions that involve the generalized Jacobians of the normal map. In particular, these two conditions are equivalent under certain structural assumptions on the generalized Jacobian matrix of the proximity operator. Leveraging second-order variational techniques and properties, we then verify that the introduced structural conditions hold for a broad class of $C^{2}$-strictly decomposable functions. Finally, it is shown that the $\mathrm{SSOSC}$ is further equivalent to the strong metric regularity of the subdifferential, the normal map, and the natural residual.
keywords:
Strong second-order sufficient condition, decomposability, second subderivative, strong metric regularity, CD-regularity, semismooth Newton method.
{AMS}
90C30, 65K05, 90C06, 90C53
1 Introduction
In this work, we consider composite-type, nonconvex, nonsmooth optimization problems of the form:
(1) $\min_{x\in\mathbb{R}^{n}}~{}\psi(x):=f(x)+\varphi(x),$
where $f:\mathbb{R}^{n}\to\mathbb{R}$ is twice continuously differentiable and $\varphi:\mathbb{R}^{n}\to(-\infty,\infty]$ is a lower semicontinuous (lsc), proper, and prox-bounded mapping. We are interested in second-order variational properties of the problem Eq. 1 and their potential implications for second-order-type methodologies.
In this context, a common strategy is to express the associated optimality conditions of Eq. 1 as a nonsmooth equation and to utilize the semismooth Newton method, [42], to solve this equation. For ease of exposition, let us suppose for a moment that $\varphi$ is weakly convex and $\tau$ is sufficiently small such that the proximity operator $\mathrm{prox}_{\tau\varphi}$ is single-valued. The natural residual equation,
(2) $F^{\tau}_{\mathrm{nat}}(x)=0,\quad F^{\tau}_{\mathrm{nat}}(x):=x-\mathrm{prox}_{\tau\varphi}(x-\tau\nabla f(x)).$
defines a popular representation of the first-order optimality conditions of Eq. 1, [16]. Another possible choice is the normal map equation, [44, 15, 38], which is given by:
(3) $F^{\tau}_{\mathrm{nor}}(z)=0,\quad F^{\tau}_{\mathrm{nor}}(z):=\nabla f(\mathrm{prox}_{\tau\varphi}(z))+{\tau^{-1}}(z-\mathrm{prox}_{\tau\varphi}(z)).$
Taking the normal map as an example, in each iteration, the semismooth Newton method generates a new step via:
(4) $z^{k+1}=z^{k}-M_{k}^{-1}F^{\tau}_{\mathrm{nor}}(z^{k}),\quad M_{k}\in\mathcal{M}^{\tau}_{\mathrm{nor}}(z^{k}),$
where $\mathcal{M}^{\tau}_{\mathrm{nor}}(z):=\{\nabla^{2}f(\mathrm{prox}_{\tau\varphi}(z))D+\frac{1}{\tau}(I-D):D\in\partial\mathrm{prox}_{\tau\varphi}(z)\}$ serves as a set of generalized Jacobian matrices of $F^{\tau}_{\mathrm{nor}}$, see, e.g., [38, 26, 30, 36] and [32, 6, 31, 51] for counterparts using the natural residual $F^{\tau}_{\mathrm{nat}}$. The superlinear convergence of the semismooth Newton steps typically requires the uniform invertibility of the matrices $M_{k}$ around the solution, i.e., the norm of the inverse of $M_{k}$ has to remain bounded by some $C$ for all sufficiently large $k$. This requirement is also related to the CD-regularity of the normal map $F^{\tau}_{\mathrm{nor}}$. One of the main goals of this paper is to address the question:
{mdframed}
[style=codebox] Can the invertibility condition, $\|M_{k}^{-1}\|\leq C$, be ensured by suitable second-order (sufficient) conditions for Eq. 1 (similar to the smooth case $\varphi\equiv 0$)?
1.1 Related Works
The semismooth Newton method was initially proposed by Qi and Sun for solving nonsmooth equations, [42]. Due to its numerical efficiency, it has been widely applied in various fields, [15, 54, 60, 6, 27, 33, 28]. The BD- and CD-regularity play central roles to guarantee the fast local superlinear convergence of the semismooth Newton method [24, 25, 41, 42, 15]. The theoretical justification of such regularity assumptions is also a key research direction in the analysis of nonsmooth Newton methods. Robinson first introduced the concept of strong regularity of the solution to a generalized equation and the strong second-order sufficient condition in [43], which includes nonlinear programming as a special case. Moreover, Robinson showed that the strong second-order sufficient condition together with the LICQ (linear independence constraint qualification) implies the strong regularity of the solution and the converse is also true for smooth nonlinear programs, see [5, Proposition 5.38] and [11, Theorem 5]. In [4], under Robinson’s constraint qualification, Bonnans and Ramírez verify that for nonlinear second-order cone programming, at a local minimizer, the strong regularity of the KKT solution mapping is equivalent to the strong second-order sufficient condition. In [52, 7], it is further shown that these notions are equivalent to the nonsingularity of Clarke’s Jacobian matrix of the KKT system for semidefinite programming. This links the CD-regularity of the stationarity equation to the strong second-order sufficient condition. Similar results are available for specific problems whose stationarity systems have explicit generalized Jacobians, which includes, e.g., Ky-Fan $k$-norm-based problems [10], matrix spectral conic programming [18], $\ell_{1}$-conic programming [29], and second-order cone programming (SOCP) [57]. Recently, a notion called variational sufficiency, [46], is shown to be equivalent to strong regularity when the set of constraints is polyhedral [46] or related to the cone of positive semidefinite matrices [55]. The framework proposed in [46] is applicable in the composite optimization setting, but direct and novel connections between the variational sufficiency condition and the uniform invertibility of the matrices $M_{k}$, $k\in\mathbb{N}$, in Eq. 4 seem mostly unavailable, due to the intrinsic hardness of calculating the generalized Jacobian of the stationarity equation. Notably, for the subspace containing derivative (SCD) mapping, BD-regularity was proven to be equivalent to the so-called SCD regularity, [17], which allows establishing the local superlinear convergence of the semismooth Newton method. The strong second-order sufficient condition is also used in [56, 55, 22] to prove superlinear convergence of certain second-order-type methods.
In the context of composite optimization, it is shown in [51, 36] that when the stationarity measure is locally smooth at the solution (which typically necessitates the strict complementarity condition to hold), the basic second-order sufficient condition can imply the uniform invertibility of the matrices $M_{k}$, $k\in\mathbb{N}$ around the solution. A similar result is established in [31, Lemma 5.4.4] by directly assuming the strict complementarity condition and $C^{2}$-fully decomposability of $\varphi$.
Without the strict complementarity condition and such local smoothness assumption, there is much less one can say about the characterization of the uniform invertibility of the generalized Jacobian of the natural residual or normal map. In [19, 50, 59, 32], $\ell_{1}$-optimization problems are considered where $\varphi$ is the $\ell_{1}$-norm. It is shown that the strong variant of the second-order sufficient condition ensures the uniform invertibility of the generalized Jacobian of the natural residual. However, it is unknown if similar results hold for a broader and more general class of functions.
It is well-known that strong metric regularity is equivalent to metric regularity for monotone operators, [23]. For convex functions, strong metric regularity is further equivalent to the uniform quadratic growth condition, [1]. For general lsc functions, strong metric regularity of the subdifferential is also equivalent to the local uniform growth condition at a local minimizer, [14]. Moreover, these conditions are equivalent to the tilt stability of the local minimizer provided that the function is prox-regular and subdifferentially continuous. Tilt stability can be also characterized via the positive definiteness of the generalized Hessian, see [14]. We are not aware of any literature concerning the equivalence between strong metric regularity of the subdifferential and strong metric regularity of the natural residual, let alone the normal map.
1.2 Contributions
We investigate a general strong second-order sufficient condition for problem Eq. 1 and its implications for uniform invertibility assumptions and the semismooth Newton method. Since the second-order differential utilized by Mordukhovich et al. in [34] lacks direct connections to the proximity operator, our analysis and formulation of the strong second-order sufficient condition will be based on the second subderivative of $\psi$, which is closely related to the directional derivative of the proximity operator and its generalized Jacobian, [39, 47]. The standard second-order sufficient conditions for Eq. 1 are given by
$\mathrm{d}^{2}\psi(\bar{x}|0)(h)\geq\sigma\|h\|^{2}\quad\forall~{}h\in\mathcal{C}(\bar{x}),$
where $\sigma>0$ and $\mathcal{C}(\bar{x}):=\{h\in\mathbb{R}^{n}:\mathrm{d}\psi(\bar{x})(h)=0\}$ is the critical cone, [47]. Mimicking the corresponding counterparts in nonlinear programming, the strong second-order sufficient condition for Eq. 1 should be expressed as a quadratic growth condition of some proper extension of $\mathrm{d}^{2}\psi(\bar{x}|0)$ on the affine hull of $\mathcal{C}(\bar{x})$. As this is not straight- forward for general $\psi$, we narrow our focus to special nonsmooth functions, whose second subderivatives have a generalized conic quadratic structure. It turns out that the class of $C^{2}$-strictly decomposable is highly “amenable” for our purposes, i.e., the second subderivative of such functions is generalized conic quadratic and, in addition, a finer characterization of the generalized Jacobian matrix of the proximity operator is available. Here, $\varphi$ is $C^{2}$-strictly decomposable at a stationary point $\bar{x}$, if $\varphi=\sigma_{\mathcal{Q}}\circ F$ is the composition of a sublinear mapping $\sigma_{\mathcal{Q}}$ and a smooth mapping $F$, a suitable constraint qualification holds at $\bar{x}$, and we have $F(\bar{x})=0$. These additional properties allow us to link the strong second-order sufficient condition to the uniform invertibility of the matrices $M_{k}$ used in Eq. 4. Surprisingly, strong metric regularity of the normal map, the subdifferential $\partial\psi$, and the natural residual can also be shown to be equivalent to the strong second-order sufficient condition. Our new results provide theoretical guarantees for the uniform invertibility assumption in the convergence analysis of the semismooth Newton method without requiring the ubiquitous strict complementarity condition, [31, 51, 36]. In summary, our contributions include:
  • We propose a strong second-order sufficient condition for nonsmooth functions whose second subderivative has a generalized conic quadratic structure, which includes the class of $C^{2}$-strictly decomposable functions as a special case.
  • We prove that the proposed strong second-order sufficient condition is equivalent to a second-order condition that is based on the generalized Jacobian of the proximity operator. This condition can be shown to be equivalent to the uniform invertibility of the matrices $M_{k}$ in Eq. 4 under certain algebraic assumptions on $\partial\mathrm{prox}_{\tau\varphi}$. Leveraging variational techniques, we verify that such algebraic conditions hold for $C^{2}$-strictly decomposable functions when a mild geometric assumption on the support set $\mathcal{Q}$, ($\varphi=\sigma_{\mathcal{Q}}\circ F$), is imposed.
  • Finally, we show that the proposed strong second-order sufficient condition is further equivalent to the strong metric regularity of the subdifferential $\partial\psi$, the natural residual $F^{\tau}_{\mathrm{nat}}$, and the normal map $F^{\tau}_{\mathrm{nor}}$.
1.3 Notation and Variational Tools
[…]
Sets. For a sequence of sets $\{C_{k}\}_{k}\subseteq\mathbb{R}^{n}$, the outer and inner limit are defined as:
${\limsup}_{k\to\infty}\,C_{k}:=\{x\in\mathbb{R}^{n}:\exists~{}\mathbb{N}\ni j_{k}\to\infty,~{}x_{j_{k}}\in C_{j_{k}},~{}x_{j_{k}}\to x\},$
${\liminf}_{k\to\infty}\,C_{k}:=\{x\in\mathbb{R}^{n}:\exists~{}x_{k}\in C_{k},~{}x_{k}\to x\}.$
Following the standard Painlevé-Kuratowski convergence for sets, we write $C_{k}\to C$ when both limits coincide. This can also be equivalently characterized by the point-wise convergence of the distance functions $\{d_{C_{k}}\}_{k}$ to $d_{C}$. For any closed set $C\subseteq\mathbb{R}^{n}$, $\Pi_{C}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ denotes the projection mapping onto $C$. If $C$ is closed and convex, then $\Pi_{C}$ is a single-valued function. In addition, if $C$ is a linear subspace, we will identify $\Pi_{C}$ with the corresponding projection matrix. We use $\iota_{C}$ and $\sigma_{C}$ to denote the indicator and support function of set $C$, respectively. The set $\mathrm{aff}(C)$ is the affine hull of $C$, $\mathrm{lin}(C)$ denotes the lineality space of $C$, and the relative interior of a convex set $C\subseteq\mathbb{R}^{n}$ is given by $\mathrm{ri}(C)$.
We also require the notion of second-order tangent sets. 1.1 is taken from [5, Definition 2.54 and 3.28].
Definition 1.1.
Assume that $C\subseteq\mathbb{R}^{n}$ is locally closed at $x\in C$. The tangent cone $T_{C}(x)$ of $C$ at $x$ is given by $T_{C}(x):=\limsup_{t\downarrow 0}{(C-x)}/{t}$. The outer second-order tangent set of $C$ at $x$ for $h$ is defined as
$T_{C}^{2}(x,h):=\limsup_{t\downarrow 0}\frac{C-x-th}{\frac{1}{2}{t^{2}}}=\left\{w:\exists~{}t_{k}\downarrow 0,\,\mathrm{dist}(x+t_{k}h+\tfrac{t_{k}^{2}}{2}w,C)=o(t^{2}_{k})\right\}$
and the inner second-order tangent set of $C$ at $x$ for $h$ is given by:
$T_{C}^{i,2}(x,h):=\liminf_{t\downarrow 0}\frac{C-x-th}{\frac{1}{2}{t^{2}}}=\left\{w:\mathrm{dist}(x+th+\tfrac{t^{2}}{2}w,C)=o(t^{2}),\,t\downarrow 0\right\}.$
For a convex set $C\subseteq\mathbb{R}^{n}$, the normal cone to $C$ at $x\in C$ is given by $N_{C}(x)=T_{C}(x)^{\circ}=\{v:\langle v,y-x\rangle\leq 0,\,\forall~{}y\in C\}$ where $\cdot^{\circ}$ denotes the polar cone operation.
Functions, epi-convergence, derivatives. The effective domain of an extended real-valued function $\theta:\mathbb{R}^{n}\to[-\infty,\infty]$ is given by $\mathrm{dom}{(\theta)}=\{x\in\mathbb{R}^{n}:\theta(x)<\infty\}$ and $\theta$ is called proper if $\theta(x)>-\infty$ and $\mathrm{dom}{(\theta)}\neq\emptyset$. The mapping $\theta$ is said to be $\rho$-weakly convex, $\rho\geq 0$, if $\theta+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. The set $\mathrm{epi}{(\theta)}:=\{(x,t)\in\mathbb{R}^{n}\times\mathbb{R}:\theta(x)\leq t\}$ denotes the epigraph of $\theta:\mathbb{R}^{n}\to(-\infty,\infty]$. A sequence of extended real-valued functions $\{\theta_{k}\}_{k}$ epi-converges to $\theta$ if $\mathrm{epi}(\theta_{k})\to\mathrm{epi}(\theta)$. Epi-convergence of $\{\theta_{k}\}_{k}$ to $\theta$ is equivalent to the condition
(5) $\left[\begin{array}[]{ll}\liminf_{k\to\infty}~{}\theta_{k}(x^{k})\,\,\geq\theta(x)&\text{for every sequence }x^{k}\to x,\\[2.15277pt] \limsup_{k\to\infty}~{}\theta_{k}(x^{k})\leq\theta(x)&\text{for some sequence }x^{k}\to x,\end{array}\right.\quad\forall~{}x.$
If the property Eq. 5 holds for fixed $x\in\mathbb{R}^{n}$, then the epi-limit of $\{\theta_{k}\}_{k}$ at $x$ exists and we write $(\operatorname*{e-lim}_{k\to\infty}\theta_{k})(x)=\theta(x)$. The lower epi-limit is given by $(\operatorname*{e-liminf}_{k\to\infty}\theta_{k})(x)=\liminf_{k\to\infty,\,\tilde{x}\to x}\theta_{k}(\tilde{x})$, cf. [47, Chapter 7].
For $\theta:\mathbb{R}^{n}\to(-\infty,\infty]$ and $x\in\mathrm{dom}(\theta)$, the regular subdifferential is defined as ^$\hat{\partial}\theta(x):=\{v\in\mathbb{R}^{n}:\theta(y)\geq\theta(x)+\langle v,y-x\rangle+o(\|y-x\|)\}$. The subdifferential of $\theta$ at $x$ is then given by $\partial\theta(x):=\limsup_{y\to x,\,\theta(y)\to\theta(x)}\hat{\partial}\theta(y)$. The second subderivative of $\theta$ at $x\in\mathrm{dom}(\theta)$ relative to $\alpha$ in the direction $h\in\mathbb{R}^{n}$ is the lower epi-limit
$\mathrm{d}^{2}\theta(x|\alpha)(h)=\liminf_{t\downarrow 0,\,\tilde{h}\to h}~{}\Delta_{t}^{2}\;\!\theta(x|\alpha)(\tilde{h}),\quad\Delta_{t}^{2}\;\!\theta(x|\alpha)(h):=\frac{\theta(x+th)-\theta(x)-t\langle\alpha,h\rangle}{\frac{1}{2}t^{2}}.$
We say that $\theta$ is twice epi-differentiable at $x$ for $\alpha\in\mathbb{R}^{n}$ if the second-order difference quotients $\Delta_{t}^{2}\;\!\theta(x|\alpha)$ epi-converge in the sense of Eq. 5.
For a mapping $F:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$, the set $\mathcal{R}(F):=\{y\in\mathbb{R}^{m}:\exists~{}x\in\mathbb{R}^{n}\,\text{with}\,y=F(x)\}$ is the range of $F$. For a locally Lipschitz continuous, vector-valued mapping $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$, the set $\partial_{B}F(x)$ denotes the Bouligand subdifferential of $F$ at $x$, i.e.,
$\partial_{B}F(x):=\{V\in\mathbb{R}^{m\times n}:x^{k}\to x,~{}\text{$F$ is differentiable at $x^{k}$},~{}DF(x^{k})\to V\}.$
Clarke’s subdifferential (generalized Jacobian) is defined as $\partial F(x)=\mathrm{conv}(\partial_{B}F(x))$. For $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$, $\mathrm{gph}(F):=\{(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{m}:y\in F(x)\}$ is the graph of $F$. Finally, we introduce the notion of metric regularity for set-valued mappings, cf. [12].
Definition 1.2.
A set-valued mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is metrically regular at $\bar{x}$ for $\bar{y}\in F(\bar{x})$, if $\mathrm{gph}(F)$ is locally closed at $(\bar{x},\bar{y})$ and there are $\kappa\geq 0$ and a neighborhood $U\times V$ of $(\bar{x},\bar{y})$ such that:
$\mathrm{dist}(x,F^{-1}(y))\leq\kappa\mathrm{dist}(y,F(x))\quad\forall~{}x\in U,\quad\forall~{}y\in V.$
If, in addition, $F^{-1}$ has a single-valued localization around $\bar{y}$ for $\bar{x}$, then $F$ is said to be strongly metrically regular at $\bar{x}$ for $\bar{y}$.
We note that a single-valued continuous mapping $F:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is strongly metrically regular at $\bar{x}$ for $F(\bar{x})$ iff there is a neighborhood $U$ of $\bar{x}$ and $\sigma>0$ such that $\sigma\|y-z\|\leq\|F(y)-F(z)\|$ for all $y,z\in U$, cf. [12, Theorem 1F.2].
Proximal properties. For a proper, lsc function $\varphi:\mathbb{R}^{n}\to(-\infty,\infty]$ and $\tau>0$, the proximity operator of $\varphi$ is defined as
$\mathrm{prox}_{\tau\varphi}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n},\quad\mathrm{prox}_{\tau\varphi}(x)\in\operatorname*{argmin}_{y\in\mathbb{R}^{n}}~{}\varphi(y)+\frac{1}{2\tau}\|x-y\|^{2},$
see [47]. The parameter $\tau$ used in the definition of $\mathrm{prox}_{\tau\varphi}$ is assumed to be positive throughout this paper. Next, we define a useful regularity concept for $\varphi$.
Definition 1.3.
We say that $\varphi$ is globally prox-regular at $\bar{x}$ for $\bar{v}\in\partial\varphi(\bar{x})$ with constant $\rho$, if $\varphi$ is locally lsc at $\bar{x}$, and there is $\epsilon>0$ such that
$\forall~{}x^{\prime}\in\mathbb{R}^{n},\quad\varphi(x^{\prime})\geq\varphi(x)+\langle v,x^{\prime}-x\rangle-\frac{\rho}{2}\|x^{\prime}-x\|^{2}$
for all $\|x-\bar{x}\|<\epsilon,\|v-\bar{v}\|<\epsilon,~{}v\in\partial\varphi(x),~{}\varphi(x)<\varphi(\bar{x})+\epsilon$.
Global prox-regularity is a mild condition and is implied by the prox-boundedness and the (standard) prox-regularity, see [47, Proposition 8.49(f) and 13.37].
Let us reconsider problem Eq. 1. The set of stationarity points of Eq. 1 is given by $\mathrm{crit}(\psi):=\{x\in\mathbb{R}^{n}:0\in\partial\psi(x)\}$. Let $\varphi$ be globally prox-regular (or prox-bounded and prox-regular) at $x$ for $v=-\nabla f(x)$ with constant $\rho$. Then, for $\tau>0$ and $\tau\rho<1$, the proximity operator $\mathrm{prox}_{\tau\varphi}$ is single-valued and Lipschitz continuous around $z=x+\tau v$, cf. [47, 37]. In this case, we can define the generalized derivates of $F^{\tau}_{\mathrm{nat}}$ and $F^{\tau}_{\mathrm{nor}}$:
$\mathcal{M}_{\mathrm{nat}}^{\tau}(x)$ :=$:=\{M:M=I-D(I-\tau\nabla^{2}f(x)),\,D\in\partial\mathrm{prox}_{\tau\varphi}(x-\tau\nabla f(x))\},$
$\mathcal{M}_{\mathrm{nor}}^{\tau}(z)$ :=$:=\{M:M=\nabla^{2}f(\mathrm{prox}_{\tau\varphi}(z))D+\tfrac{1}{\tau}(I-D),\,D\in\partial\mathrm{prox}_{\tau\varphi}(z)\}.$
By [8], we have $\partial F^{\tau}_{\mathrm{nor}}(z)h=\mathcal{M}_{\mathrm{nor}}^{\tau}(z)h$ for all $h\in\mathbb{R}^{n}$. In addition, if $x\in\mathrm{crit}(\psi)$, then it follows $F^{\tau}_{\mathrm{nat}}(x)=0$, $F^{\tau}_{\mathrm{nor}}(z)=0$, and $x=\mathrm{prox}_{\tau\varphi}(z)$, see, e.g., [36, Lemma 2.1].
Convex geometry. Finally, we discuss several concepts from convex geometry that will be used in Subsection 3.2. The face and exposed face of a convex set are defined as follows, cf. [20, 58]:
Definition 1.4.
For a convex set $C\subseteq\mathbb{R}^{n}$, a face $F$ of $C$ is a subset of $C$ with the property that $\lambda y+(1-\lambda)z=x\in F$ with $y,z\in C$ and $\lambda\in(0,1)$ implies $[y,z]\subseteq F$. An exposed face $F_{\mathrm{ex}}$ is a subset of $C$ which can be expressed as $F_{\mathrm{ex}}=H\cap C$, where $H$ is a supporting hyperplane of $C$. Equivalently, $F_{\mathrm{ex}}$ is an exposed face of $C$ if there exists $x\neq 0$ such that $F_{\mathrm{ex}}=\{y\in C:\langle x,y\rangle=\sigma_{C}(x)\}$. Let $\mathcal{F}_{\mathrm{ex}}$ denote the set of all exposed faces of $C$. The smallest exposed face of $C$ containing $x\in C$ is denoted by $G_{\mathrm{ex}}(x):=\bigcap\,\{F\in\mathcal{F}_{\mathrm{ex}}:x\in F\}$.
Noticing $\partial\sigma_{C}(x)=\{y\in C:\langle x,y\rangle=\sigma_{C}(x)\}$, each exposed face $F_{\mathrm{ex}}$ of the set $C$ can be represented by $\partial\sigma_{C}(x)$ for an appropriate point $x\neq 0$. Next, we collect several well-known properties of the smallest exposed face $G_{\mathrm{ex}}$, cf. [58, Lemma 4.6] and [48, Lemma 2.2.2]. We provide a self-contained proof of 1.5 in Appendix A.
Lemma 1.5.
Let $C\subseteq\mathbb{R}^{n}$ be a convex set and let $x\in C$ be given.
  1. (i)
    If $y\in\mathrm{ri}(G_{\mathrm{ex}}(x))$, then we have $N_{C}(x)=N_{C}(y)$ and $T_{C}(x)=T_{C}(y)$.
  2. (ii)
    It holds that $G_{\mathrm{ex}}(x)=\partial\sigma_{C}(y)$ for all $y\in\mathrm{ri}(N_{C}(x))$.
1.4 Organization
This paper is organized as follows. In Section 2, we introduce the strong second-order sufficient optimality condition ($\mathrm{SSOSC}$) and a related second-order-type condition that is based on the generalized derivatives of the normal map $F^{\tau}_{\mathrm{nor}}$. We then formulate structural assumptions on $\partial\mathrm{prox}_{\tau\varphi}$ to ensure equivalence between those two optimality conditions. In Section 3, we analyze the second-order variational properties of strictly decomposable functions and we verify that the postulated structural assumptions on $\partial\mathrm{prox}_{\tau\varphi}$ hold if $\varphi$ is $C^{2}$-strictly decomposable under a mild, additional assumption on the support set of the decomposition. In Section 4, we prove equivalence between the $\mathrm{SSOSC}$ and other variational concepts, including the strong metric regularity of the subdifferential $\partial\psi$, the normal map, and the natural residual, and several other second-order-type conditions.
2 Connecting the $\mathrm{SSOSC}$ and Generalized Jacobians
In this section, we formally define the $\mathrm{SSOSC}$ and discuss its equivalence to a second-order condition which can ensure the uniform invertibility of the matrices $M_{k}$, $k\in\mathbb{N}$. Specifically, we propose two algebraic assumptions on the generalized Jacobian matrix of the proximity operator that can guarantee such equivalence.
The strong second-order condition given in [35, 34] fully characterizes the strong metric regularity of the subdifferential of the objective function by using the second-order subdifferential based on Mordukhovich’s coderivative. However, a drawback of this characterization is the missing link to generalized Jacobians of the stationarity equation (i.e., the underlying KKT system, the natural residual, or the normal map equation) and to the semismooth Newton method. This is of course understandable as the calculation of the generalized Jacobian is nontrivial in general. As a result, connections between the $\mathrm{SSOSC}$ and the generalized Jacobian of the stationarity equation are only known for problems whose stationarity equation has explicit, closed-form generalized Jacobians, [52, 57, 18, 29].
Since the second subderivative and the derivative of the proximity operator are closely connected, [39], we plan to use the second subderivative rather than the second-order differential to formulate the $\mathrm{SSOSC}$. According to [47, Theorem 13.24(c)], the standard second-order sufficient condition takes the form $\mathrm{d}^{2}\psi(\bar{x}|0)(h)>0$ for all $h\neq 0$. By the homogeneity and lower semicontinuity of the second subderivative, this condition can be equivalently expressed as
(6) $\mathrm{d}^{2}\psi(\bar{x}|0)(h)\geq\sigma\|h\|^{2}\quad\forall~{}h\in\mathrm{dom}(\mathrm{d}^{2}\psi(\bar{x}|0))$
for some $\sigma>0$. As in the nonlinear programming case, [4, 52], we expect the $\mathrm{SSOSC}$ to strenghten Eq. 6 using a proper extension of $\mathrm{d}^{2}\psi(\bar{x}|0)$ on the affine hull of its domain. Such a generalization is not always easy, since, in contrast to the specific applications considered in the literature, the second subderivative of $\varphi$ can be too complex to be extended naturally. Hence, we narrow our focus on a special class of functions whose second subderivative provides a natural extension to the affine hull of the critical cone.
{assumption}
Let $\bar{x}\in\mathrm{crit}{(\psi)}$, $\bar{v}=-\nabla f(\bar{x})$, and $\bar{z}=\bar{x}+\tau\bar{v}$ be given.
  1. (A.1)
    The function $\varphi$ is globally prox-regular at $\bar{x}$ for $\bar{v}$ with constant $\rho$ and $\tau\rho<1$.
  2. (A.2)
    The function $\varphi$ is twice epi-differentiable at $\bar{x}$ for $\bar{v}$ and its second subderivative $\mathrm{d}^{2}\varphi(\bar{x}|\bar{v})$ has a generalized conic quadratic form, i.e.,
    $\mathrm{d}^{2}\varphi(\bar{x}|\bar{v})(h)=\langle h,Qh\rangle+\iota_{S}(h)\quad\forall~{}h\in\mathbb{R}^{n},$
    where $Q\in\mathbb{R}^{n\times n}$ is some symmetric matrix with $\mathcal{R}(Q)\subseteq\mathrm{aff}(S)$ and $S\subseteq\mathbb{R}^{n}$ is a closed convex cone.
For simplicity, we will often use the notation $\mathsf{A}:=\mathrm{aff}(S)$. We further note that the condition $\mathcal{R}(Q)\subseteq\mathsf{A}$ causes no loss of generality, since we can replace $Q$ by $\Pi_{\mathsf{A}}Q\Pi_{\mathsf{A}}$.
(A.1) is a default condition to ensure the single-valuedness and Lipschitz continuity of the proximity operator around $\bar{z}$ (as mentioned in Subsection 1.3). We assume (A.1) throughout this paper. Under assumption (A.2) and applying [47, Exercise 13.18], the second subderivative $\mathrm{d}^{2}\psi(\bar{x}|0)$ satisfies
(7) $\mathrm{d}^{2}\psi(\bar{x}|0)(h)=\langle h,[\nabla^{2}f(\bar{x})+Q]h\rangle+\iota_{S}(h),$
and thus, it also has a generalized conic quadratic form. In this case, an extension of $\mathrm{d}^{2}\psi(\bar{x}|0)$ from its domain to the affine hull of its domain can be obtained naturally and directly and we can define the following strong second-order condition.
Definition 2.1 (Strong Second-order Sufficient Condition).
Let $\mathrm{d}^{2}\psi(\bar{x}|0)$ have the form Eq. 7 for a closed convex cone $S$ and $Q\in\mathbb{S}^{n}$ with $\mathcal{R}(Q)\subseteq\mathrm{aff}(S)$. The strong second-order sufficient condition $(\mathrm{SSOSC})$ holds at $\bar{x}$ if there is $\sigma>0$ such that:
(8) $\langle h,[\nabla^{2}f(\bar{x})+Q]h\rangle\geq\sigma\|h\|^{2}\quad\forall~{}h\in\mathrm{aff}(S).$
Motivated by the results in [36], we consider a second condition that involves the generalized Jacobian of the proximity operator:
(9) $D\nabla^{2}f(\bar{x})D+{\tau}^{-1}D(I-D)\succeq\sigma D^{2}\quad\forall~{}D\in\partial\mathrm{prox}_{\tau\varphi}(\bar{z}).$
As shown in [36], condition Eq. 9 guarantees CD-regularity of the normal map $F^{\tau}_{\mathrm{nor}}$ at $\bar{z}$. To illustrate this fact, let $M\in\partial F^{\tau}_{\mathrm{nor}}(\bar{z})$ and $h\in\mathbb{R}^{n}$ with $Mh=0$ be arbitrary. Due to $\partial F^{\tau}_{\mathrm{nor}}(\bar{z})h=\mathcal{M}_{\mathrm{nor}}^{\tau}(\bar{z})h$, there is $D\in\partial\mathrm{prox}_{\tau\varphi}(\bar{z})$ such that $Mh=[\nabla^{2}f(\bar{x})D+{\tau}^{-1}(I-D)]h$. By Eq. 9, it then follows $Dh=0$ and we have $0=[\nabla^{2}f(\bar{x})-{\tau}^{-1}I]Dh+{\tau}^{-1}h={\tau^{-1}}h$. This verifies invertibility of $M$ and the desired CD-regularity.
We are interested in the relationship between Eq. 8 and Eq. 9. Clearly, a proper characterization of the generalized Jacobians $\partial\mathrm{prox}_{\tau\varphi}$ is integral to establish a connection between Eq. 8 and Eq. 9. The most direct way to verify the equivalence of Eq. 8 and Eq. 9 is to calculate $\partial\mathrm{prox}_{\tau\varphi}$, $Q$, and $\mathrm{aff}(S)$ explicitly. Indeed, such a direct computational verification has been the prevalent proof strategy in applications and has been used successfully in, e.g., nonlinear semidefinite programming [52, 7], second order cone programs [4, 56, 57], and Ky Fan $k$-norm cone programming [9]. In the following, we propose several abstract conditions on $\partial\mathrm{prox}_{\tau\varphi}$ under which Eq. 8 is equivalent to Eq. 9. Utilizing variational techniques, we then show that these conditions hold for a broad class of strictly decomposable problems.
2.1 A First Structural Condition for $\partial\mathrm{prox}_{\tau\varphi}$
By assumption (A.1) and [37, Proposition 2.3], it holds that $0\preceq D\preceq\frac{1}{1-\tau\rho}I$ for all $D\in\partial\mathrm{prox}_{\tau\varphi}(\bar{z})$. In addition, the matrix $I+\tau Q$ is positive definite. We start with a simple algebraic observation.
Lemma 2.2.
Let $D\in\mathbb{S}^{n}$ be given with $0\preceq D\preceq\frac{1}{1-\tau\rho}I$. Then, $D$ can be written as $D=\Pi_{\mathcal{R}(D)}(I+\tau A)^{-1}\Pi_{\mathcal{R}(D)}$, where $A\in\mathbb{S}^{n}$ satisfies $A+\rho I\succeq 0$ and $\mathcal{R}(A)\subseteq\mathcal{R}(D)$.
Proof.
Let $D=P\Lambda P^{\top}$ be an eigendecomposition of $D$ with $PP^{\top}=P^{\top}P=I$ and eigenvalues $\operatorname{diag}(\lambda)=\Lambda$. Let us assume $\lambda_{i}>0$ for $i=1,\dots,r$ and $\lambda_{i}=0$ for all $i=r+1,\dots,n$ and let us further write $P=[U,V]$ where $U\in\mathbb{R}^{n\times r}$ and $V\in\mathbb{R}^{n\times(n-r)}$. Defining $\Lambda_{1}:=\mathrm{diag}(\lambda_{1},\dots,\lambda_{r})$ and
$A:=\frac{1}{\tau}P\begin{pmatrix}\Lambda_{1}^{-1}-I&0\\ 0&0\end{pmatrix}P^{\top}=\frac{1}{\tau}U(\Lambda_{1}^{-1}-I)U^{\top},$
it follows $\Pi_{\mathcal{R}(D)}(I+\tau A)^{-1}\Pi_{\mathcal{R}(D)}=U\Lambda_{1}U^{\top}=D$, $\mathcal{R}(A)\subseteq\mathcal{R}(D)$, $\Lambda_{1}^{-1}\succeq(1-\tau\rho)I$, and $A+\rho I\succeq 0$.
2.2 motivates a first structural condition for the generalized Jacobians:
(P.1) for every Dproxτφ(z¯), we $\begin{array}[]{l}\textit{for every $D\in\partial\mathrm{prox}_{\tau\varphi}(\bar{z})$, we have:}\\[4.30554pt] \hskip 4.30554ptD=\Pi_{\mathcal{R}(D)}(I+\tau A)^{-1}\Pi_{\mathcal{R}(D)},\;\;\mathcal{R}(A)\subseteq\mathcal{R}(D)\subseteq\mathsf{A},\;\;A\succeq\Pi_{\mathcal{R}(D)}Q\Pi_{\mathcal{R}(D)},\end{array}$
where $\mathsf{A}=\mathrm{aff}(S)$. The positive definiteness of the matrix $I+\tau Q$ ensures that $(I+\tau A)^{-1}$ is well-defined and a bijective mapping from $\mathcal{R}(D)$ to $\mathcal{R}(D)$. We now show that, under condition Eq. P.1, the strong second-order sufficient condition Eq. 8 implies Eq. 9.
Proposition 2.3.
Assume (A.1), (A.2), and Eq. P.1. Then, Eq. 8 implies Eq. 9.
Proof.
By Eq. P.1 and setting $\mathcal{S}:=\mathcal{R}(D)$, the condition $D\nabla^{2}f(\bar{x})D+\frac{1}{\tau}D(I-D)\succeq\sigma D^{2}$ can be written as
$\Pi_{\mathcal{S}}(I+\tau A)^{-1}\Pi_{\mathcal{S}}\nabla^{2}f(\bar{x})\Pi_{\mathcal{S}}(I+\tau A)^{-1}\Pi_{\mathcal{S}}$
$\hskip 17.22217pt+\Pi_{\mathcal{S}}(I+\tau A)^{-1}A(I+\tau A)^{-1}\Pi_{\mathcal{S}}\succeq\sigma\Pi_{\mathcal{S}}(I+\tau A)^{-2}\Pi_{\mathcal{S}},$
which is equivalent to
$\langle(I+\tau A)^{-1}h,(\Pi_{\mathcal{S}}\nabla^{2}f(\bar{x})\Pi_{\mathcal{S}}+A)(I+\tau A)^{-1}h\rangle\geq\sigma\|(I+\tau A)^{-1}h\|^{2}\quad\forall~{}h\in\mathcal{S}.$
Notice that since $\mathcal{R}(A)\subseteq\mathcal{S}$, $(I+\tau A)^{-1}$ is a bijective mapping from $\mathcal{S}$ to $\mathcal{S}$. Hence, replacing $(I+\tau A)^{-1}h$ by $h$, the last condition is further equivalent to:
$\langle h,(\Pi_{\mathcal{S}}\nabla^{2}f(\bar{x})\Pi_{\mathcal{S}}+A)h\rangle\geq\sigma\|h\|^{2}\quad\forall~{}h\in\mathcal{S}.$
Using Eq. P.1 and Eq. 8, it follows $\Pi_{\mathcal{S}}\nabla^{2}f(\bar{x})\Pi_{\mathcal{S}}+A\succeq\Pi_{\mathcal{S}}\nabla^{2}f(\bar{x})\Pi_{\mathcal{S}}+\Pi_{\mathcal{S}}Q\Pi_{\mathcal{S}}\succeq\sigma\Pi_{\mathcal{S}}^{2}$ which concludes the proof.
Mimicking the proof of 2.3, we can show that the condition
(P.2) $\Pi_{\mathsf{A}}(I+\tau Q)^{-1}\Pi_{\mathsf{A}}\in\partial\mathrm{prox}_{\tau\varphi}(\bar{z}).$
ensures the implication “Eq. 9 $\implies$ Eq. 8”. We will omit an explicit verification here.
Proposition 2.4.
Assume (A.1), (A.2), and Eq. P.2. Then Eq. 9 implies Eq. 8.
2.2 A Weaker Form of Eq. P.1
We now consider a weaker variant of the condition Eq. P.1 that only involves the Bouligand subdifferential $\partial_{B}\mathrm{prox}_{\tau\varphi}(\bar{z})$:
(P.3) for every DBproxτφ(z¯), we $\begin{array}[]{l}\textit{for every $D\in\partial_{B}\mathrm{prox}_{\tau\varphi}(\bar{z})$, we have:}\\[4.30554pt] \hskip 4.30554ptD=\Pi_{\mathcal{R}(D)}(I+\tau A)^{-1}\Pi_{\mathcal{R}(D)},\quad\mathcal{R}(A)\subseteq\mathcal{R}(D)\subseteq\mathsf{A},\quad A\succeq\Pi_{\mathcal{R}(D)}Q\Pi_{\mathcal{R}(D)}.\end{array}$
Our goal is to prove that Eq. P.3 already implies the stronger condition Eq. P.1. We first present a result on the range of the sum of positive semidefinite matrices.
Proposition 2.5.
Let $\{A_{i}\}_{i=1,\dots,r}$ be a family of symmetric positive semidefinite matrices and let $a_{i}>0$, $i=1,\dots,r$, be given. Then, $\mathcal{R}(\sum_{i=1}^{r}a_{i}A_{i})=\sum_{i=1}^{r}\mathcal{R}(A_{i})$ where the sum of subspaces is defined as $\sum_{i=1}^{r}S_{i}:=\{\sum_{i=1}^{r}x_{i}:x_{i}\in S_{i},i=1,\dots,r\}$.
Proof.
We note that for a symmetric matrix $A$, it follows $\ker(A)=\mathcal{R}(A)^{\perp}$ and for two positive semidefinite matrices $A$, $B$, we have $\ker(A)\cap\ker(B)=\ker(A+B)$. Therefore by [2, Propositions 6.26 and 6.34], it holds that
$\mathcal{R}\left({\sum}_{i=1}^{r}a_{i}A_{i}\right)=\ker\left({\sum}_{i=1}^{r}a_{i}A_{i}\right)^{\perp}=\left({\bigcap}_{i=1}^{r}\ker(A_{i})\right)^{\perp}={\sum}_{i=1}^{r}\mathcal{R}(A_{i}).$
Next, based on the representation $D=\Pi_{\mathcal{R}(D)}(I+\tau A)^{-1}\Pi_{\mathcal{R}(D)}$ in Eq. P.3, we discuss the order of positive semidefinite matrices under projections.
Lemma 2.6.
Let $L\subseteq\mathbb{R}^{n}$ be a linear subspace of $\mathbb{R}^{n}$ and let $B\in\mathbb{S}^{n}$ be given with $I+B\succ 0$. Then, it holds that $\Pi_{L}(I+\Pi_{L}B\Pi_{L})^{-1}\Pi_{L}\preceq(I+B)^{-1}$.
Proof.
Without loss of generality, let us assume $L=\mathrm{span}\{e_{1},\dots,e_{r}\}$ and let us write $B$ in block form:
$B=\begin{pmatrix}B_{1}&B_{2}\\ B_{2}^{\top}&B_{3}\end{pmatrix},\quad B_{1}\in\mathbb{R}^{r\times r},\quad B_{2}\in\mathbb{R}^{r\times n-r},\quad B_{3}\in\mathbb{R}^{n-r\times n-r}.$
By the block inverse formula [21, Equation (0.8.5.6)] and using the Schur complement $C=(I+B_{3})-B_{2}^{\top}(I+B_{1})^{-1}B_{2}$ for $I+B$, we see that:
$(I+B)^{-1}=\begin{pmatrix}(I+B_{1})^{-1}+(I+B_{1})^{-1}B_{2}C^{-1}B_{2}^{\top}(I+B_{1})^{-1}&-(I+B_{1})^{-1}B_{2}C^{-1}\\ -C^{-1}B_{2}^{\top}(I+B_{1})^{-1}&C^{-1}\end{pmatrix},$
$\Pi_{L}(I+\Pi_{L}B\Pi_{L})^{-1}\Pi_{L}=\begin{pmatrix}(I+B_{1})^{-1}&0\\ 0&0\end{pmatrix}.$
Notice that $C^{-1}$ is positive definite as principle submatrix of the positive definite matrix $(I+B)^{-1}$. Moreover, we have
$(I+B)^{-1}-\Pi_{L}(I+\Pi_{L}B\Pi_{L})^{-1}\Pi_{L}$
$\hskip 17.22217pt=\begin{pmatrix}(I+B_{1})^{-1}B_{2}C^{-1}B_{2}^{\top}(I+B_{1})^{-1}&-(I+B_{1})^{-1}B_{2}C^{-1}\\ -C^{-1}B_{2}^{\top}(I+B_{1})^{-1}&C^{-1}\end{pmatrix}=:E.$
To conclude, it suffices to show that $E$ is positive semidefinite. But taking arbitrary $h=(h_{1},h_{2})\in\mathbb{R}^{n}$, it follows $h^{\top}Eh=\|C^{-\frac{1}{2}}B_{2}^{\top}(I+B_{1})^{-1}h_{1}-C^{-\frac{1}{2}}h_{2}\|^{2}\geq 0$.
Proposition 2.7.
Let $B\in\mathbb{S}^{n}$ with $I+\tau B\succ 0$ be given and let $\{S_{i}\}_{i=1,\dots,r}$ be a family of linear subspaces of $\mathbb{R}^{n}$. For $i=1,\dots,r$ consider $P_{i}=\Pi_{S_{i}}(I+\tau A_{i})^{-1}\Pi_{S_{i}}$ with $A_{i}\in\mathbb{S}^{n}$, $\mathcal{R}(A_{i})\subseteq S_{i}$, and $A_{i}\succeq\Pi_{S_{i}}B\Pi_{S_{i}}$ and set $P=\sum_{i=1}^{r}a_{i}P_{i}$ where $a_{i}\geq 0$ and $\sum_{i=1}^{r}a_{i}=1$. Then, the matrix $P$ can be written as $P=\Pi_{\mathcal{R}(P)}(I+\tau A)^{-1}\Pi_{\mathcal{R}(P)}$ with $\mathcal{R}(A)\subseteq\mathcal{R}(P)$ and $A\succeq\Pi_{\mathcal{R}(P)}B\Pi_{\mathcal{R}(P)}$. In particular, Eq. P.3 implies Eq. P.1.
Proof.
We first note that by combining $A_{i}\succeq\Pi_{S_{i}}B\Pi_{S_{i}}$ and $I+\tau B\succ 0$, we can infer that each $I+\tau A_{i}$ is positive definite. We prove 2.7 by induction. The base case $r=1$ is trivial. Let us now consider $r=2$. In the case $a_{1}=0$ or $a_{2}=0$ nothing needs to be shown, so we can assume $a_{1},a_{2}>0$. We then have $\mathcal{R}(P)=S_{1}+S_{2}$ and setting $\tilde{B}:=\Pi_{S_{1}+S_{2}}B\Pi_{S_{1}+S_{2}}$, it follows:
$A_{1}\succeq\Pi_{S_{1}}\tilde{B}\Pi_{S_{1}}\quad\text{and}\quad A_{2}\succeq\Pi_{S_{2}}\tilde{B}\Pi_{S_{2}}.$
Utilizing 2.6, we obtain
$\Pi_{S_{1}}(I+\tau A_{1})^{-1}\Pi_{S_{1}}\preceq\Pi_{S_{1}}(I+\tau\Pi_{S_{1}}\tilde{B}\Pi_{S_{1}})^{-1}\Pi_{S_{1}}\preceq(I+\tau\tilde{B})^{-1}$
and $\Pi_{S_{2}}(I+\tau A_{2})^{-1}\Pi_{S_{2}}\preceq(I+\tau\tilde{B})^{-1}$ which implies $P\preceq(I+\tau\tilde{B})^{-1}$. By 2.2, we can write $P$ as $P=\Pi_{S_{1}+S_{2}}(I+\tau A)^{-1}\Pi_{S_{1}+S_{2}}$ for some $A$ with $\mathcal{R}(A)\subseteq\mathcal{R}(P)=S_{1}+S_{2}$. This yields $\Pi_{S_{1}+S_{2}}(I+\tau A)^{-1}\Pi_{S_{1}+S_{2}}\preceq(I+\tau\tilde{B})^{-1}$. Since both $\mathcal{R}(A)$ and $\mathcal{R}(\tilde{B})$ are subsets of $S_{1}+S_{2}$, this condition is true if and only if:
$(I+\tau A)^{-1}\preceq(I+\tau\tilde{B})^{-1}~\quad\iff\quad A\succeq\tilde{B},$
which proves the case $r=2$. Now suppose the conclusion is true for $11,\dots,r-$ 1. There is no loss of generality, if we assume $a_{i}>0$ for all $i$. Let us set $\tilde{P}=\frac{1}{1-a_{r}}\sum_{i=1}^{r-1}a_{i}P_{i}$. Using induction for $\tilde{P}$ and applying the last steps for $P=(1-a_{r})\tilde{P}+a_{r}P_{r}$, we can conclude the first part of the proof. The remaining assertion in 2.7 follows from $\partial\mathrm{prox}_{\tau\varphi}(\bar{z})=\mathrm{conv}(\partial_{B}\mathrm{prox}_{\tau\varphi}(\bar{z}))$ and the Carathéodory theorem, cf. [47].
2.3 Second-order Variational Formulations for Eq. P.2 and Eq. P.3
Due to the close relationship between the generalized Jacobian of the proximity operator $\mathrm{prox}_{\tau\varphi}$ and the second subderivative of $\varphi$, we now plan to convert Eq. P.2 and Eq. P.3 into corresponding second-order variational conditions. The main tool is [37, Lemma 3.13], which allows connecting the limit of $D\mathrm{prox}_{\tau\varphi}$ and the epi-limit of the second subderivative of $\varphi$ with each other.
Proposition 2.8.
Let (A.1) and (A.2) be satisfied. Assume further that for all $\mathrm{gph}(\partial\varphi)\ni(x^{k},v^{k})\to(\bar{x},\bar{v})$ such that $\varphi$ is twice epi-differentiable at $x^{k}$ for $v^{k}$ and $\mathrm{d}^{2}\varphi(x^{k}|v^{k})$ is generalized quadratic, it holds that:
(V.1) $({\operatorname*{e-liminf}}_{k\to\infty}\,\mathrm{d}^{2}\varphi(x^{k}|v^{k}))(w)\geq\langle w,Qw\rangle+\iota_{\mathrm{aff}(S)}(w)\quad\forall~{}w\in\mathbb{R}^{n}.$
Then, condition Eq. P.3 is satisfied.
Proof.
Let $D\in\partial_{B}\mathrm{prox}_{\tau\varphi}(\bar{z})$ be arbitrary. Due to [37, Proposition 2.3], we have $0\preceq D\preceq\frac{1}{1-\tau\rho}I$ and applying 2.2, it holds that $D=\Pi_{\mathcal{R}(D)}(I+\tau A)^{-1}\Pi_{\mathcal{R}(D)}$ with $A\in\mathbb{S}^{n}$, $A+\rho I\succeq 0$, and $\mathcal{R}(A)\subseteq\mathcal{R}(D)$. By definition, there is $z^{k}\to\bar{z}$ such that $\mathrm{prox}_{\tau\varphi}$ is differentiable at $z^{k}$ and $D\mathrm{prox}_{\tau\varphi}(z^{k})\to D$. Let us define $x^{k}:=\mathrm{prox}_{\tau\varphi}(z^{k})$ and $v^{k}:=\frac{1}{\tau}(z^{k}-x^{k})$. Then, it follows $z^{k}=x^{k}+\tau v^{k}$ and $v^{k}\in\partial\varphi(x^{k})$. Applying [37, Lemma 3.13], the condition $D\mathrm{prox}_{\tau\varphi}(z^{k})\to D$ implies that $\varphi$ is twice epi-differentiable at $x^{k}$ for $v^{k}$, $\mathrm{d}^{2}\varphi(x^{k}|v^{k})$ is generalized quadratic (for all $k$), and we have
$({\operatorname*{e-lim}}_{k\to\infty}\,\mathrm{d}^{2}\varphi(x^{k}|v^{k}))(w)=\langle w,Aw\rangle+\iota_{\mathcal{R}(D)}(w)\quad\forall~{}w\in\mathbb{R}^{n}.$
By assumption, we know that:
$\langle w,Aw\rangle+\iota_{\mathcal{R}(D)}(w)\geq\langle w,Qw\rangle+\iota_{\mathrm{aff}(S)}(w)\quad\forall~{}w\in\mathbb{R}^{n}.$
This yields $\Pi_{\mathcal{R}(D)}Q\Pi_{\mathcal{R}(D)}\preceq A$ and $\mathcal{R}(A)\subseteq\mathcal{R}(D)\subseteq\mathrm{aff}(S)$ which proves Eq. P.3.
Proposition 2.9.
Let (A.1) and (A.2) hold. Assume further that there exists a sequence $\mathrm{gph}(\partial\varphi)\ni(x^{k},v^{k})\to(\bar{x},\bar{v})$ such that $\varphi$ is twice epi-differentiable at $x^{k}$ for $v^{k}$, $\mathrm{d}^{2}\varphi(x^{k}|v^{k})$ is generalized quadratic, and we have
(V.2) $({\operatorname*{e-lim}}_{k\to\infty}\,\mathrm{d}^{2}\varphi(x^{k}|v^{k}))(w)=\langle w,Qw\rangle+\iota_{\mathrm{aff}(S)}(w)\quad\forall~{}w\in\mathbb{R}^{n}.$
Then, condition Eq. P.2 is satisfied.
Proof.
This is a direct consequence of [37, Lemma 3.13].
3 Second-order Variational Properties of $C^{2}$-Strictly Decomposable Functions
In this section, we introduce the concept of $C^{2}$-strict decomposability. Strictly decomposable functions satisfy condition (A.2) and we show that Eq. V.1 and Eq. V.2 will hold under additional mild geometric assumptions.
Definition 3.1 (Decomposable functions).
A function $\varphi:\mathbb{R}^{n}\rightarrow(-\infty,\infty]$ is called $C^{\ell}$-decomposable, $\ell\in\mathbb{N}$, at a point $\bar{x}\in\mathrm{dom}(\varphi)$, if there is an open neighborhood $U$ of $\bar{x}$ such that
$\varphi(x)=\varphi(\bar{x})+\varphi_{d}(F(x))\quad\forall~{}x\in U,$
and the functions $\varphi_{d}$ and $F$ satisfy:
  1. (i)
    $F:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is $\ell$-times continuously differentiable on $U$ and $F(\bar{x})=0$.
  2. (ii)
    The mapping $\varphi_{d}:\mathbb{R}^{m}\rightarrow(-\infty,\infty]$ is convex, lsc, positively homogeneous, and proper, or equivalently, $\varphi_{d}=\sigma_{\mathcal{Q}}$ for some closed convex set $\mathcal{Q}$.
  3. (iii)
    Robinson’s constraint qualification holds at $\bar{x}$ :
    $0\in\operatorname{int}\left\{F(\bar{x})+DF(\bar{x})\mathbb{R}^{n}-\mathrm{dom}(\varphi_{d})\right\}=\operatorname{int}\left\{DF(\bar{x})\mathbb{R}^{n}-\mathrm{dom}(\varphi_{d})\right\}.$
We say that $\varphi$ is $C^{\ell}$-strictly decomposable at $\bar{x}$ for $\bar{\lambda}$ if $\varphi$ is $C^{\ell}$-decomposable at $\bar{x}$ and if, in addition, the strict condition
(10) $DF(\bar{x})\mathbb{R}^{n}-N_{\mathcal{Q}}(\bar{\lambda})=\mathbb{R}^{m}$
is satisfied at $\bar{x}$ for some $\bar{\lambda}\in\mathcal{Q}$. We say that $\varphi$ is $C^{\ell}$-fully decomposable at $\bar{x}$ if the strict condition is replaced by the nondegeneracy condition $DF(\bar{x})\mathbb{R}^{n}+\operatorname{lin}(N_{\mathcal{Q}}(\bar{\lambda}))=\mathbb{R}^{m}$.
The concept of decomposability was first proposed by Shapiro [49] and serves as a functional analogue to cone reducible sets. Examples of $C^{2}$-strictly decomposable functions include, e.g., polyhedral and piecewise-linear functions, the total variation and the group-lasso norm, and the Ky-Fan $k$-norm. For more examples and background, we refer to [49, 31, 37].
By [37, Proposition 3.7], if $\varphi$ is $C^{2}$-strictly decomposable at $\bar{x}$ for $\bar{\lambda}\in\mathcal{Q}$, then $\varphi$ is properly twice epi-differentiable at $\bar{x}$ for $\bar{v}=DF(\bar{x})^{\top}\bar{\lambda}\in\partial\varphi(\bar{x})$ with
(11) $\mathrm{d}^{2}\varphi(\bar{x}|\bar{v})(w)=\langle\bar{\lambda},D^{2}F(\bar{x})[w,w]\rangle+\iota_{DF(\bar{x})^{-1}N_{\mathcal{Q}}(\bar{\lambda})}(w)\quad\forall~{}w\in\mathbb{R}^{n}.$
Here, $\bar{\lambda}\in\mathcal{Q}$ is uniquely determined by $\bar{v}\in\partial\varphi(\bar{x})$ thanks to [37, Proposition 3.6]. Defining $H(x,\lambda):=\sum_{i=1}^{m}\lambda_{i}\nabla^{2}F_{i}(x)$, assumption (A.2) is then clearly satisfied with $S=DF(\bar{x})^{-1}N_{\mathcal{Q}}(\bar{\lambda})$ and $Q=\Pi_{\mathsf{A}}H(\bar{x},\bar{\lambda})\Pi_{\mathsf{A}}$. In the following, we verify that the conditions Eq. V.1 and Eq. V.2 hold for $C^{2}$-strictly decomposable functions. As in [37] and as mentioned, this requires an additional geometric assumption on the support set $\mathcal{Q}$.
3.1 Verifying Eq. V.1 and the $\mathcal{OTP}$-Property
Definition 3.2.
We say that a closed convex set $\mathcal{Q}$ has the outer second-order tangent path ($\mathcal{OTP}$-)property at $\bar{\lambda}\in\mathcal{Q}$, if there is a constant $M>0$ and a neighborhood $V$ of $\bar{\lambda}$ such that for all $\lambda\in\mathcal{Q}\cap V$, it holds that:
  1. (i)
    We can select a linear subspace $L(\lambda)\subseteq\mathrm{lin}(T_{\mathcal{Q}}(\lambda))$ such that $L(\lambda)\to\mathrm{lin}(T_{\mathcal{Q}}(\bar{\lambda}))$ as $\lambda\to\bar{\lambda}$;
  2. (ii)
    For all $p\in L(\lambda)$ with $\|p\|=1$, we have $\mathrm{dist}(0,T^{2}_{\mathcal{Q}}(\lambda,p))<M$.
If this holds for all $\bar{\lambda}\in\mathcal{Q}$, then we say that $\mathcal{Q}$ has the $\mathcal{OTP}$-property.
The $\mathcal{OTP}$-property is a weaker variant of the uniform second-order tangent path ($\mathcal{UTP}$-)property introduced in [37, Definition 3.16]. Specifically, compared to the $\mathcal{UTP}$-property, uniform length of the second-order path along the direction $p\in L(\lambda)$ is not required and we only need the curvature of the second-order outer tangent set to be uniformly bounded rather than the inner one. Since we only need to control the second subderivative $\mathrm{d}^{2}\varphi$, which is defined as limit of second-order difference quotients, uniformness of the tangent paths is not necessary. In addition, since $\varphi$ is assumed to be twice epi-differentiable at $\bar{x}$ for $\bar{v}$ in Eq. V.1 and Eq. V.2, the inner second-order tangent set can be replaced by the outer one (as illustrated in the proof of 3.6).
Since the $\mathcal{OTP}$-property is implied by the $\mathcal{UTP}$-property, all sets having the $\mathcal{UTP}$-property also have $\mathcal{OTP}$-property. In particular, this includes $C^{2}$-pointed and $C^{2}$-cone reducible sets, cf. [37].
Next, we construct two counterexamples to illustrate that both conditions in 3.2 are necessary for the implication “Eq. 8 $\implies$ Eq. 9”.
(a) Visualization for Example 3.3
(b) Visualization for Example 3.4
Figure 1: Illustration of the sets and constructions in Example 3.3 and Example 3.4.
Example 3.3 (Necessity of condition $\mathrm{(i)}$).
We consider:
$C_{1}=\{(0,0,x_{3}):x_{3}\in[-1,1]\},\quad C_{2}=\{(x_{1},x_{2},0):x_{1}^{2}+(x_{2}-1)^{2}=1,\,x_{1}\geq 0\},$
$C=\overline{\mathrm{conv}}(C_{1}\cup C_{2}),\quad\varphi_{d}(x)=\sigma_{C}(x),\quad F=I,\quad\varphi=\varphi_{d},\quad\text{$f$ free},$
and $\bar{x}=\bar{\lambda}=(0,0,0)^{\top}$. Then, it holds that $\bar{\lambda}\in\partial\varphi(0)=C$ and it follows $N_{C}(\bar{\lambda})=\{(x_{1},x_{2},0)^{\top}:x_{1},x_{2}\leq 0\}$. By construction, $\varphi$ is $C^{2}$-fully decomposable at $\bar{x}$ with decomposition pair $(\varphi_{d},F)$. Hence, using Eq. 11 with $\bar{v}=DF(\bar{x})^{\top}\bar{\lambda}=0$, we have
$\mathrm{d}^{2}\varphi(\bar{x}|0)(h)=\langle h,H(\bar{x},\bar{\lambda})h\rangle+\iota_{N_{C}(\bar{\lambda})}(h)=\iota_{N_{C}(\bar{\lambda})}(h)\quad\forall~{}h\in\mathbb{R}^{n}$
and the strong second-order sufficient condition Eq. 8 is equivalent to:
(12) $\langle h,\nabla^{2}f(\bar{x})h\rangle\geq\sigma\|h\|^{2}\quad\forall~{}h\in\mathrm{aff}(N_{C}(\bar{\lambda}))=\mathrm{span}\{(1,0,0)^{\top},(0,1,0)^{\top}\}.$
Next, let us define $x^{k}:=\frac{1}{k}(\frac{1}{k},-\sqrt{1-1/k^{2}},0)^{\top}$ and $\lambda^{k}:=(\frac{1}{k},1-\sqrt{1-{1}/{k^{2}}},0)^{\top}$. We first claim that $\lambda^{k}\in\partial\varphi(x^{k})=\partial\varphi_{d}(x^{k})=\{\lambda\in C:\langle\lambda,x^{k}\rangle=\varphi_{d}(x^{k})\}$, which requires to compute $\varphi_{d}$. Let $x\in K:=\{x\in\mathbb{R}^{3}:x_{1}>0,\,x_{2}<0,\,\text{and}\,|x_{3}|<\sqrt{x_{1}^{2}+x_{2}^{2}}+x_{2}\}$ be given. Applying [20, Theorem 3.3.2], we can calculate:
$\sigma_{C}(x)=\max\{\sigma_{C_{1}}(x),\sigma_{C_{2}}(x)\}=\max\{|x_{3}|,\sqrt{x_{1}^{2}+x_{2}^{2}}+x_{2}\}=\sqrt{x_{1}^{2}+x_{2}^{2}}+x_{2}.$
This expression is valid for all $x\in K$ and hence, $\varphi$ is twice continuously differentiable at such $x$. The gradient and Hessian of $\varphi$ at $x$ are given by:
$\nabla\varphi(x)=\frac{1}{n(x)}\begin{pmatrix}{x_{1}}\\ n(x)+{x_{2}}\\ 0\end{pmatrix},\quad\nabla^{2}\varphi(x)=\frac{1}{n(x)^{3}}\begin{pmatrix}{n(x)^{2}}-{x_{1}^{2}}&-{x_{1}x_{2}}&0\\ -{x_{1}x_{2}}&{n(x)^{2}}-{x_{2}^{2}}&0\\ 0&0&0\end{pmatrix},$
where $n(x):=\sqrt{x_{1}^{2}+x_{2}^{2}}$. We then obtain $x^{k}\in K$, $\lambda^{k}\in C_{2}$, and $\varphi_{d}(x^{k})=\langle\lambda^{k},x^{k}\rangle$ for all $k$, which verifies $\lambda^{k}\in\partial\varphi(x^{k})$. In addition, using the implicit function theorem, $\mathrm{prox}_{\tau\varphi}$ is differentiable at $z^{k}=x^{k}+\tau\lambda^{k}$ with $D\mathrm{prox}_{\tau\varphi}(z^{k})=(I+\tau\nabla^{2}\varphi(x^{k}))^{-1}$. (Notice that $\lambda^{k}\in\partial\varphi(x^{k})$ implies $x^{k}=\mathrm{prox}_{\tau\varphi}(z^{k})$). It holds that $z^{k}\to(0,0,0)^{\top}$ and
$D\mathrm{prox}_{\tau\varphi}(z^{k})=(I+\tau\nabla^{2}\varphi(x^{k}))^{-1}\to\mathrm{diag}(0,1,1)=:\bar{D}\quad\text{as}\;k\to\infty.$
Thus, the second-order condition Eq. 9 for $\bar{D}\in\partial\mathrm{prox}_{\tau\varphi}(0)$ requires:
$\langle h,\nabla^{2}f(\bar{x})h\rangle\geq\sigma\|h\|^{2}\quad\forall~{}h\in\mathrm{span}\{(0,1,0)^{\top},(0,0,1)^{\top}\}.$
Of course, this is not implied by condition Eq. 12. Furthermore, Eq. P.1 and Eq. V.1 fail to hold. In this example, condition $\mathrm{(i)}$ in the definition of the $\mathcal{OTP}$-property is not satisfied. We actually have $\mathrm{lin}(T_{C}(\bar{\lambda}))=\{(0,0,t):t\in\mathbb{R}\}$ but $\mathrm{lin}(T_{C}(\lambda^{k}))\to\{(t,0,0):t\in\mathbb{R}\}$.
Example 3.4 (Necessity of condition $\mathrm{(ii)}$).
We consider [37, Example 3.21]:
$C=\{x\in\mathbb{R}^{2}:x_{2}\geq\tfrac{2}{3}|x_{1}|^{\frac{3}{2}}\},\quad F=I,\quad\varphi(x)=\varphi_{d}(x)=\sigma_{C}(x),\quad\text{$f$ free}$
with $\bar{x}=\bar{\lambda}=(0,0)^{\top}$. Then, it holds that $\bar{\lambda}\in\partial\varphi(0)$,
$N_{C}(\bar{\lambda})=\{(0,t):t\leq 0\},\quad\text{and}\quad\mathrm{lin}(T_{C}(\bar{\lambda}))=\{(t,0):t\in\mathbb{R}\}.$
In addition, we have $\mathrm{lin}(T_{C}(\lambda))\to\mathrm{lin}(T_{C}(\bar{\lambda}))$ for $C\backslash\mathrm{ri}(C)\ni\lambda\to\bar{\lambda}$. Therefore, the first condition in the definition of the $\mathcal{OTP}$-property is satisfied. However, the second condition does not hold since it can be shown that the outer second-order tangent sets at $\bar{\lambda}$ for the directions $(1,0)$ and $(-1,0)$ are both empty. Similar to Example 3.3 and by construction of $\varphi$, it follows $\mathrm{d}^{2}\varphi(\bar{x}|\bar{\lambda})(h)=\iota_{N_{C}(\bar{\lambda})}(h)$ for all $h$. Hence, condition Eq. 8 is again equivalent to:
$\langle h,\nabla^{2}f(\bar{x})h\rangle\geq\sigma\|h\|^{2}\quad\forall~{}h\in\mathrm{aff}(N_{C}(\bar{\lambda}))=\mathrm{span}\{(0,1)^{\top}\}.$
Let us set $x^{k}=(\frac{1}{k^{3}},-\frac{1}{k})^{\top}$ and $\lambda^{k}=(\frac{1}{k^{4}},\frac{1}{k^{6}})^{\top}$. As calculated in [37, Example 3.21], it holds that $\lambda^{k}\in\partial\varphi(x^{k})$ and
$\nabla^{2}\varphi(x^{k})=2\begin{pmatrix}\frac{1}{k}&-\frac{1}{k^{3}}\\ -\frac{1}{k^{3}}&\frac{1}{k^{5}}\end{pmatrix}\to 0.$
Setting $z^{k}=x^{k}+\tau\lambda^{k}$, we have $\mathrm{prox}_{\tau\varphi}(z^{k})=x^{k}$, $z^{k}\to\bar{z}=\bar{x}$ and we further obtain:
$D\mathrm{prox}_{\tau\varphi}(z^{k})=(I+\tau\nabla^{2}\varphi(x^{k}))^{-1}\to I\in\partial\mathrm{prox}_{\tau\varphi}(\bar{z}).$
Thus, Eq. 9 requires $\nabla^{2}f(\bar{x})\succeq\sigma I$. Consequently, condition Eq. 9 is stronger than Eq. 8. Moreover, condition Eq. P.1 fails, which implies that Eq. V.1 also fails.
Recall that condition (A.1) is assumed throughout this paper. We now list the main assumptions to ensure Eq. V.1 and “Eq. 8 $\implies$ Eq. 9”:
Assumption 3.5.
We consider the conditions:
  1. (B.1)
    $\varphi$ is $C^{2}$-strictly decomposable at $\bar{x}$ for $\bar{\lambda}$ with decomposition pair $(\varphi_{d}\equiv\sigma_{\mathcal{Q}},F)$.
  2. (B.2)
    In addition to (B.1), suppose that $\mathcal{Q}$ has the $\mathcal{OTP}$-property at $\bar{\lambda}$.
As in [37], our first task is to exploit the postulated geometric properties of the support set $\mathcal{Q}$ in the estimation of the second subderivative of $\varphi$.
The following technical lemma is our main tool.
Lemma 3.6.
Suppose that $\varphi$ has the decomposition $\varphi=\sigma_{\mathcal{Q}}\circ F$ around $x\in\mathrm{dom}(\varphi)$ and let $\lambda\in\partial\sigma_{\mathcal{Q}}(F(x))$, $p\in T_{\mathcal{Q}}(\lambda)$, $w\in\mathbb{R}^{n}$ be given with $\langle p,DF(x)w\rangle>0$ and $\langle p,F(x)\rangle=0$. Furthermore, assume that $\mathrm{dist}(0,T^{2}_{\mathcal{Q}}(\lambda,p))<M$ and $\varphi$ is twice epi-differentiable at $x$ for $DF(x)^{\top}\lambda$. Then, it holds that
$\mathrm{d}^{2}\varphi(x|DF(x)^{\top}\lambda)(w)\geq\frac{|\langle p,DF(x)w\rangle|^{2}}{M\|F(x)\|}+\langle w,H(x,\lambda)w\rangle,$
where the right-hand side is understood as $\infty$ if $F(x)=0$.
Proof.
The proof of 3.6 is inspired by the proof of [37, Lemma 3.23]. The condition $\mathrm{dist}(0,T^{2}_{\mathcal{Q}}(\lambda,p))<M$ implies that there is a sequence $\{\xi^{k}\}_{k}$ such that
$\xi^{k}=\lambda+s_{k}p+\frac{1}{2}s_{k}^{2}r^{k},\quad\xi^{k}\in\mathcal{Q},\quad s_{k}\downarrow 0,\quad\|r^{k}\|\leq M.$
The twice epi-differentiability of $\varphi$ at $x$ for $DF(x)^{\top}\lambda$ says that for any choice of $t_{k}\downarrow 0$, we can find $w^{k}\to w$ such that
$\mathrm{d}^{2}\varphi(x|DF(x)^{\top}\lambda)(w)=\lim_{k\to\infty}\Delta_{t_{k}}^{2}\;\!\varphi(x|DF(x)^{\top}\lambda)(w^{k}).$
We now select $t_{k}$ as follows:
$t_{k}=\begin{cases}s_{k}^{2}&\text{if }F(x)=0,\\ \frac{Ms_{k}\|F(x)\|}{\langle p,DF(x)w\rangle}&\text{if }F(x)\neq 0.\end{cases}$
In particular, in the case $F(x)\neq 0$, this choice implies
(13) $s_{k}=\frac{t_{k}\langle p,DF(x)w\rangle}{M\|F(x)\|}.$
Noticing $\partial\sigma_{\mathcal{Q}}(F(x))=\{\mu\in\mathcal{Q}:\langle\mu,F(x)\rangle=\sigma_{\mathcal{Q}}(F(x))\}$, we now have:
$\Delta_{t_{k}}^{2}\;\!\varphi(x|DF(x)^{\top}\lambda)(w^{k})=\frac{\sigma_{\mathcal{Q}}(F(x+t_{k}w^{k}))-\sigma_{\mathcal{Q}}(F(x))-t_{k}\langle DF(x)^{\top}\lambda,w^{k}\rangle}{\frac{1}{2}t_{k}^{2}}$
$\hskip 25.83325pt\geq\frac{\langle\xi^{k},F(x+t_{k}w^{k})\rangle-\langle\lambda,F(x)\rangle-t_{k}\langle DF(x)^{\top}\lambda,w^{k}\rangle}{\frac{1}{2}t_{k}^{2}}$
$\hskip 25.83325pt=\frac{\langle s_{k}p+\frac{1}{2}s_{k}^{2}r^{k},F(x+t_{k}w^{k})\rangle}{\frac{1}{2}t_{k}^{2}}+\frac{\langle\lambda,F(x+t_{k}w^{k})-F(x)-t_{k}DF(x)w^{k}\rangle}{\frac{1}{2}t_{k}^{2}}$
$\hskip 25.83325pt=\frac{\langle s_{k}p+\frac{1}{2}s_{k}^{2}r^{k},F(x)+t_{k}DF(x)w^{k}+\frac{1}{2}t_{k}^{2}E(t_{k},w^{k})\rangle}{\frac{1}{2}t_{k}^{2}}+\langle\lambda,E(t_{k},w^{k})\rangle,$
where $E(t_{k},w^{k}):=\int_{0}^{1}D^{2}F(x+st_{k}w^{k})[w^{k},w^{k}]\,\mathrm{d}s\in\mathbb{R}^{m}$. Utilizing the assumption $\langle p,F(x)\rangle=0$, we further obtain
$\langle s_{k}p+\tfrac{1}{2}s_{k}^{2}r^{k},F(x)+t_{k}DF(x)w^{k}+\tfrac{1}{2}t_{k}^{2}E(t_{k},w^{k})\rangle$
$=s_{k}t_{k}\langle p,DF(x)w^{k}\rangle+\tfrac{1}{2}s_{k}^{2}\langle r^{k},F(x)\rangle+\tfrac{1}{2}s_{k}^{2}t_{k}\langle r^{k},DF(x)w^{k}\rangle+o(t_{k}^{2})$
Hence, if $F(x)=0$, it follows:
$\langle s_{k}p+\tfrac{1}{2}s_{k}^{2}r^{k},F(x)+t_{k}DF(x)w^{k}+\tfrac{1}{2}t_{k}^{2}E(t_{k},w^{k})\rangle=t_{k}^{\frac{3}{2}}\langle p,DF(x)w^{k}\rangle+\mathcal{O}(t_{k}^{2})$
and due to $\langle p,DF(x)w^{k}\rangle\to\langle p,DF(x)w\rangle>0$, we can infer
$\Delta_{t_{k}}^{2}\;\!\varphi(x|DF(x)^{\top}\lambda)(w^{k})$ $\geq\frac{t_{k}^{\frac{3}{2}}\langle p,DF(x)w^{k}\rangle+\mathcal{O}(t_{k}^{2})}{\frac{1}{2}t_{k}^{2}}+\langle\lambda,E(t_{k},w^{k})\rangle\to\infty.$
This proves the conclusion when $F(x)=0$. Next, let us consider the case $F(x)\neq 0$. Then, by Eq. 13, we have:
$\langle s_{k}p+\tfrac{1}{2}s_{k}^{2}r^{k},F(x)+t_{k}DF(x)w^{k}+\tfrac{1}{2}t_{k}^{2}E(t_{k},w^{k})\rangle$
$=s_{k}t_{k}\langle p,DF(x)w^{k}\rangle+\frac{1}{2}s_{k}^{2}\langle r^{k},F(x)\rangle+o(t_{k}^{2})$
$\geq s_{k}[t_{k}\langle p,DF(x)w^{k}\rangle-\frac{1}{2}s_{k}\|r^{k}\|\|F(x)\|]+o(t_{k}^{2})$
$=\frac{t_{k}^{2}}{M\|F(x)\|}\left[\langle p,DF(x)w^{k}\rangle\langle p,DF(x)w\rangle-\frac{\|r^{k}\|}{2M}\langle p,DF(x)w\rangle^{2}\right]+o(t_{k}^{2}).$
Therefore, taking the limit $k\to\infty$, it holds that
$\lim_{k\to\infty}\frac{\langle s_{k}p+\frac{1}{2}s_{k}^{2}r^{k},F(x)+t_{k}DF(x)w^{k}+\frac{1}{2}t_{k}^{2}E(t_{k},w^{k})\rangle}{\frac{1}{2}t_{k}^{2}}\geq\frac{|\langle p,DF(x)w\rangle|^{2}}{M\|F(x)\|}.$
Combining the previous steps and calculations, this yields
$\mathrm{d}^{2}\varphi(x|DF(x)^{\top}\lambda)(w)$ $\geq\frac{|\langle p,DF(x)w\rangle|^{2}}{M\|F(x)\|}+\langle w,H(x,\lambda)w\rangle,$
which finishes the proof.
Thanks to 3.6, we can now establish the key result of this subsection.
theorem 3.7.
Let (B.1) be satisfied and suppose there is a sequence $\mathrm{gph}(\partial\varphi)\ni(x^{k},v^{k})\to(\bar{x},\bar{v})$ such that $\varphi$ is twice epi-differentiable at $x^{k}$ for $v^{k}$ for all $k\in\mathbb{N}$. Assume further that there exist a constant $M>0$ and $\{\lambda^{k}\}_{k}$ with $v^{k}=DF(x^{k})^{\top}\lambda^{k}$ and $\lambda^{k}\in\partial\sigma_{\mathcal{Q}}(F(x^{k}))$ such that the following conditions hold:
  1. (i)
    There are linear subspaces $L_{k}\subseteq\mathrm{lin}(T_{\mathcal{Q}}(\lambda^{k}))$ such that $L_{k}\to\mathrm{lin}(T_{\mathcal{Q}}(\bar{\lambda}))$,
  2. (ii)
    For all $p\in L_{k}$ with $\|p\|=1$, we have $\mathrm{dist}(0,T^{2}_{\mathcal{Q}}(\lambda^{k},p))<M$.
Then, it holds that:
(14) ${\operatorname*{e-liminf}}_{k\to\infty}\,\mathrm{d}^{2}\varphi(x^{k}|DF(x^{k})^{\top}\lambda^{k})\geq\langle\cdot,H(\bar{x},\bar{\lambda})\cdot\rangle+\iota_{\mathrm{aff}(DF(\bar{x})^{-1}N_{\mathcal{Q}}(\bar{\lambda}))}.$
In particular, if (B.1) and (B.2) hold, then condition Eq. V.1 is satisfied.
Proof.
Using [3, Corollary 4.3] and [47, Proposition 13.20], we have $\mathrm{d}^{2}\varphi(x^{k}|v^{k})\geq\langle\cdot,H(x^{k},\lambda^{k})\cdot\rangle$ and applying [37, Proposition 3.6], it follows $\lambda^{k}\to\bar{\lambda}$. Thus, taking the epi-liminf and noticing that $\langle\cdot,H(x^{k},\lambda^{k})\cdot\rangle$ converges continuously, we obtain
(15) ${\operatorname*{e-liminf}}_{k\to\infty}\,\mathrm{d}^{2}\varphi(x^{k}|v^{k})\geq\langle\cdot,H(\bar{x},\bar{\lambda})\cdot\rangle.$
Applying [37, Lemma 3.22], we know $\mathrm{aff}(DF(\bar{x})^{-1}N_{\mathcal{Q}}(\bar{\lambda}))=DF(\bar{x})^{-1}\mathrm{aff}(N_{\mathcal{Q}}(\bar{\lambda}))$. Let us now take $w\notin DF(\bar{x})^{-1}(\mathrm{aff}(N_{\mathcal{Q}}(\bar{\lambda})))$, which implies $\|\Pi_{\mathrm{lin}(T_{\mathcal{Q}}(\bar{\lambda}))}(DF(\bar{x})w)\|=2\epsilon$ for some $\epsilon>0$. By the definition of the epi-liminf, we may find $w^{k}\to w$ such that
$({\operatorname*{e-liminf}}_{k\to\infty}\,\mathrm{d}^{2}\varphi(x^{k}|v^{k}))(w)={\liminf}_{k\to\infty}\,\mathrm{d}^{2}\varphi(x^{k}|v^{k})(w^{k}).$
Notice that the set convergence $L_{k}\to\mathrm{lin}(T_{\mathcal{Q}}(\bar{\lambda}))$ implies $\Pi_{L_{k}}\to\Pi_{\mathrm{lin}(T_{\mathcal{Q}}(\bar{\lambda})}$. Hence, by continuity, we may assume $\|\Pi_{L_{k}}(DF(x^{k})w^{k})\|>\epsilon$ for all $k\in\mathbb{N}$. Next, let us set $p^{k}={\Pi_{L_{k}}(DF(x^{k})w^{k})}/{\|\Pi_{L_{k}}(DF(x^{k})w^{k})\|}$. Then, due to $p^{k}\in L_{k}\subseteq\mathrm{lin}(T_{\mathcal{Q}}(\lambda^{k}))$ and $F(x^{k})\in\partial\sigma_{\mathcal{Q}}^{*}(\lambda^{k})=N_{\mathcal{Q}}(\lambda^{k})$, it holds that $\langle p^{k},F(x^{k})\rangle=0$ and $\langle p^{k},DF(x^{k})w^{k}\rangle=\|\Pi_{L_{k}}(DF(x^{k})w^{k})\|>\epsilon$. Applying 3.6 to $x^{k}$, $\lambda^{k}$, $w^{k}$ and $p^{k}$, we have
$\mathrm{d}^{2}\varphi(x^{k}|DF(x^{k})^{\top}\lambda^{k})(w^{k})\geq\frac{\|\Pi_{L_{k}}(DF(x^{k})w^{k})\|^{2}}{2M\|F(x^{k})\|}+\langle w^{k},H(x^{k},\lambda^{k})w^{k}\rangle$
for all $k$. Consequently, taking the limit and using $F(x^{k})\to F(\bar{x})=0$, we obtain:
(16) $({\operatorname*{e-liminf}}_{k\to\infty}\,\mathrm{d}^{2}\varphi(x^{k}|v^{k}))(w)=\infty.$
Therefore, Eq. 14 follows from Eq. 15 and Eq. 16. Under assumption (B.1), it holds that $\partial\varphi(x^{k})=DF(x^{k})^{\top}\partial\sigma_{\mathcal{Q}}(F(x^{k}))$ (for all $k$ sufficiently large), cf. [37, Proposition 3.5]. Thus, the required properties of $\{\lambda^{k}\}_{k}$ and the conditions (i) and (ii) are ensured under (B.1) and (B.2). Condition Eq. V.1 then follows from Eq. 14.
3.2 Verifying Eq. V.2
As illustrated in 3.7, the geometric requirements (B.2) on $\mathcal{Q}$ allowed us to establish an appropriate lower bound for the second subderivative of $\varphi$. In order to verify Eq. V.2, the generalized quadratic structure of the second subderivative – as stated in 2.9 – will play a more important role. Since the function $\varphi$ is often $C^{2}$-strictly decomposable everywhere [31, 37], we plan to use [31, Lemma 5.3.27] and [37, Theorem 3.11] to connect the generalized quadratic structure of $\mathrm{d}^{2}\varphi$ to the strict complementarity condition.
Moreover, to verify Eq. V.2, we only need to find a specific sequence $\{(x^{k},v^{k})\}_{k}$ rather than controlling all possible sequences as in Eq. V.1. Therefore, the conditions formulated in the $\mathcal{OTP}$-property are not required to hold on a neighborhood of $\bar{\lambda}$.
Assumption 3.8.
We consider the conditions:
  1. (B.3)
    For all $x$ on a neighborhood of $\bar{x}$, the strict complementarity condition $v\in\mathrm{ri}(\partial\varphi(x))$ implies that $\mathrm{prox}_{\tau\varphi}$ is differentiable at $x+\tau v$.
  2. (B.4)
    There exists $\bar{M}>0$ such that for all $v\in\mathrm{lin}(T_{\mathcal{Q}}(\bar{\lambda}))$ with $\|v\|=1$, we have $\mathrm{dist}(0,T_{\mathcal{Q}}^{2}(\bar{\lambda},v))<\bar{M}$.
Remark 3.9.
We note that assumption (B.3) holds if $\varphi$ is $C^{2}$-strictly decomposable on some neighborhood of $\bar{x}$, cf. [37, Theorem 3.11].
As the normal cone $N_{\mathcal{Q}}(\bar{\lambda})$ and the smallest exposed face $G_{\mathrm{ex}}({\bar{\lambda}})$ of $\mathcal{Q}$ containing $\bar{\lambda}$ will be utilized frequently in this subsection, we introduce the notations $\mathcal{N}:=N_{\mathcal{Q}}(\bar{\lambda})$ and $\mathcal{G}:=G_{\mathrm{ex}}({\bar{\lambda}})$. By [37, Proposition 3.4(ii)], we know that the strict constraint qualification $DF(\bar{x})\mathbb{R}^{n}-\mathcal{N}=\mathbb{R}^{m}$ (see Eq. 10) is stable on a neighborhood of $\bar{x}$.
Proposition 3.10.
Let condition (B.1) hold. Then there is a neighborhood $U$ of $\bar{x}$ such that for all $x\in U$ it holds that $DF(x)\mathbb{R}^{n}-\mathcal{N}=\mathbb{R}^{m}$.
To verify Eq. V.2, we need to select a proper sequence $(x^{k},v^{k})\to(\bar{x},\bar{v})$ to satisfy the conditions listed in 2.9. We start with the choice of $\{x^{k}\}_{k}$.
Proposition 3.11.
Let (B.1) hold. There exists $\{x^{k}\}_{k}$ with $x^{k}\to\bar{x}$, $F(x^{k})\in\mathrm{ri}(\mathcal{N})$, $\partial\varphi(x^{k})=DF(x^{k})^{\top}\mathcal{G}$ (for all $k$), and $DF(x^{k})^{-1}\mathrm{aff}(\mathcal{N})\to DF(\bar{x})^{-1}\mathrm{aff}(\mathcal{N})$.
Proof.
This follows from [37, Proposition 3.9]. The representation $\partial\varphi(x^{k})=DF(x^{k})^{\top}\mathcal{G}$ is due to [37, Proposition 3.5] and 1.5 (ii).
We now choose an appropriate vector $\lambda^{k}$ and construct $v^{k}$ via $v^{k}=DF(x^{k})^{\top}\lambda^{k}$. We require the condition $\lambda^{k}\in\mathrm{ri}(\mathcal{G})$ to apply (B.3). Moreover, (B.2) should hold for $\lambda^{k}$ allowing us to use 3.7. We specify $\lambda^{k}$ in the following proposition.
Proposition 3.12.
Let $\bar{\lambda}\in\mathcal{Q}$ be given and let $\mathcal{G}:=G_{\mathrm{ex}}({\bar{\lambda}})$ denote the smallest exposed face of $\mathcal{Q}$ containing $\bar{\lambda}$. Furthermore, suppose that $T^{2}_{\mathcal{Q}}(\bar{\lambda},p)\neq\emptyset$ for all $p\in\mathrm{lin}(T_{\mathcal{Q}}(\bar{\lambda}))$. Then, there exists a sequence $\{\lambda^{k}\}_{k}\subseteq\mathrm{ri}(\mathcal{G})$ with $\lambda^{k}\to\bar{\lambda}$ such that $\mathrm{dist}(0,T_{\mathcal{Q}}^{2}(\lambda^{k},p))\leq 2\mathrm{dist}(0,T_{\mathcal{Q}}^{2}(\bar{\lambda},p))$ for all $p\in\mathrm{lin}(T_{\mathcal{Q}}(\bar{\lambda}))$ and $k\in\mathbb{N}$.
Proof.
We fix $\tilde{\lambda}\in\mathrm{ri}(\mathcal{G})$. Take $\nu\in(0,1)$ and let us set $\lambda=\nu\bar{\lambda}+(1-\nu)\tilde{\lambda}$. By [45, Theorem 6.1], it follows $\lambda\in\mathrm{ri}(\mathcal{G})$. By 1.5 (i), we have $T_{\mathcal{Q}}(\lambda)=T_{\mathcal{Q}}(\bar{\lambda})$. Next, let $p\in\mathrm{lin}(T_{\mathcal{Q}}(\bar{\lambda}))$ be arbitrary. By assumption, there is a sequence $\{\xi^{n}\}_{n}$ with $\xi^{n}=\bar{\lambda}+t_{n}p+\frac{1}{2}t_{n}^{2}r^{n}\in\mathcal{Q}$, $t_{n}\downarrow 0$, $r^{n}\to r$, and $\|r\|=\mathrm{dist}(0,T_{\mathcal{Q}}^{2}(\bar{\lambda},p))$. (Since $T^{2}_{\mathcal{Q}}(\bar{\lambda},p)$ is closed as set limit, the projection of 00 onto $T^{2}_{\mathcal{Q}}(\bar{\lambda},p)$ exists). This immediately yields:
$\zeta^{n}:=\nu\xi^{n}+(1-\nu)\tilde{\lambda}=\lambda+\nu t_{n}p+\tfrac{1}{2}(\nu t_{n})^{2}({r^{n}}/{\nu})\in\mathcal{Q}.$
By definition, this then implies $\frac{r}{\nu}\in T_{C}^{2}(\lambda,p)$ and $\mathrm{dist}(0,T^{2}_{\mathcal{Q}}(\lambda,p))\leq\frac{1}{\nu}\mathrm{dist}(0,T_{\mathcal{Q}}^{2}(\bar{\lambda},p))$. Therefore, it suffices to choose $\lambda^{k}:=\nu_{k}\bar{\lambda}+(1-\nu_{k})\tilde{\lambda}$ with $\nu_{k}\geq\frac{1}{2}$ and $\nu_{k}\to 1$.
Our next task is to give an exact expression of $\mathrm{d}^{2}\varphi(x^{k}|DF(x^{k})^{\top}\lambda^{k})$ (on the linear subspace $DF(x^{k})^{-1}\mathrm{aff}(\mathcal{N})$) to calculate the epi-limit.
Proposition 3.13.
Let (B.1) be satisfied. Then, there exists a neighborhood $U$ of $\bar{x}$ such that for all $x\in U$ with $F(x)\in\mathrm{ri}(\mathcal{N})$ and $\lambda\in\mathrm{ri}(\mathcal{G})$, it holds that
$\mathrm{d}^{2}\varphi(x|DF(x)^{\top}\lambda)(w)=\langle w,H(x,\lambda)w\rangle,\quad\forall~{}w\in DF(x)^{-1}\mathrm{aff}(\mathcal{N}).$
Proof.
By 3.10, there is a neighborhood $U$ of $\bar{x}$ such that $DF(x)\mathbb{R}^{n}-\mathcal{N}=\mathbb{R}^{m}$ holds for all $x\in U$. By [3, Corollary 4.3], we again have $\mathrm{d}^{2}\varphi(x|DF(x)^{\top}\lambda)\geq\langle\cdot,H(x,\lambda)\cdot\rangle$. Hence, it suffices to prove the converse inequality. Now assume $w\in DF(x)^{-1}\mathrm{aff}(\mathcal{N})$ and define $K:=\{x\in\mathbb{R}^{n}:F(x)\in\mathcal{N}\}$. Applying [5, Corollary 2.91] and $F(x)\in\mathrm{ri}(\mathcal{N})$, it follows $T_{K}(x)=DF(x)^{-1}T_{\mathcal{N}}(F(x))=DF(x)^{-1}\mathrm{aff}(\mathcal{N})$, see [20, Section A.5.3]. Thus, we can infer $w\in T_{K}(x)$ and due to $\lambda\in\mathrm{ri}(\mathcal{G})$ and 1.5 (i), it holds that $N_{\mathcal{Q}}(\lambda)=\mathcal{N}$. Hence, there exist $t_{k}\downarrow 0$ and $w^{k}\to w$ such that $F(x+t_{k}w^{k})\in\mathcal{N}=N_{\mathcal{Q}}(\lambda)$. By the definition of the second subderivative, this yields:
$\mathrm{d}^{2}\varphi(x|DF(x)^{\top}\lambda)(w)\leq\lim_{k\to\infty}\Delta_{t_{k}}^{2}\;\!(\sigma_{\mathcal{Q}}\circ F)(x|DF(x)^{\top}\lambda)(w^{k})$
$=\lim_{k\to\infty}\frac{\langle\lambda,F(x+t_{k}w^{k})-F(x)-t_{k}DF(x)w^{k}\rangle}{\frac{1}{2}t_{k}^{2}}=\langle w,H(x,\lambda),w\rangle.$
We now combine the different propositions to show one of our main results.
theorem 3.14.
Assume (B.1), (B.3) and (B.4). Then, condition Eq. V.2 is satisfied.
Proof.
Let $\{x^{k}\}_{k}$ and $\{\lambda^{k}\}_{k}$ be given as in 3.10 and 3.11. The two conditions $\lambda^{k}\in\mathrm{ri}(\mathcal{G})$ and $\partial\varphi(x^{k})=DF(x^{k})^{\top}\mathcal{G}$ imply $v^{k}:=DF(x^{k})^{\top}\lambda^{k}\in\mathrm{ri}(\partial\varphi(x^{k}))$ by [45, Theorem 6.6]. Thus, by (B.3) and [37, Proposition 2.2 and 2.3], $\mathrm{prox}_{\tau\varphi}$ is differentiable at $z^{k}=x^{k}+\tau v^{k}$ with $0\preceq D\mathrm{prox}_{\tau\varphi}(z^{k})\preceq\frac{1}{1-\tau\rho}I$. Taking a subsequence if necessary, we may assume $D\mathrm{prox}_{\tau\varphi}(z^{k})\to R$ for some $0\preceq R\preceq\frac{1}{1-\tau\rho}I$. Applying [37, Proposition 2.2], we know that $\varphi$ is twice epi-differentiable at $x^{k}$ for $v^{k}$, $\mathrm{d}^{2}\varphi(x^{k}|v^{k})$ is generalized quadratic and converges epigraphically. Noticing $T_{\mathcal{Q}}(\lambda^{k})=T_{\mathcal{Q}}(\bar{\lambda})$ and using 3.12, (B.4), and 3.7, we have
(17) ${\operatorname*{e-liminf}}_{k\to\infty}\,\mathrm{d}^{2}\varphi(x^{k}|v^{k})\geq\langle\cdot,H(\bar{x},\bar{\lambda})\cdot\rangle+\iota_{\mathrm{aff}(DF(\bar{x})^{-1}\mathcal{N})}.$
As before, by [37, Lemma 3.22], it holds that $\mathrm{aff}(DF(\bar{x})^{-1}\mathcal{N})=DF(\bar{x})^{-1}\mathrm{aff}(\mathcal{N})$. Let $w\in DF(\bar{x})^{-1}\mathrm{aff}(\mathcal{N})$ be arbitrary. By 3.11, we can choose $w^{k}\to w$ with $w^{k}\in DF(x^{k})^{-1}\mathrm{aff}(\mathcal{N})$. Applying 3.13, we obtain
(18) $({\operatorname*{e-liminf}}_{k\to\infty}\,\mathrm{d}^{2}\varphi(x^{k}|v^{k}))(w)\leq{\lim}_{k\to\infty}\,\mathrm{d}^{2}\varphi(x^{k}|v^{k})(w^{k})=\langle w,H(\bar{x},\bar{\lambda})w\rangle.$
Condition Eq. V.2 then follows from Eq. 17 and Eq. 18 and by noticing that the sequence $\{\mathrm{d}^{2}\varphi(x^{k}|v^{k})\}_{k}$ epi-converges.
4 Equivalence between Other Conditions
In this section, we investigate the equivalence between the $\mathrm{SSOSC}$ and other variational concepts for Eq. 1. Specifically, we show that the strong second-order sufficient condition, the strong metric regularity of the subdifferential, the normal-map, the natural residual, and the uniform invertibility of the generalized derivatives in $\mathcal{M}_{\mathrm{nat}}^{\tau}(\bar{x})$ and $\mathcal{M}_{\mathrm{nor}}^{\tau}(\bar{z})$ are all equivalent under the setting studied in Section 3. Our aim is to prove the following theorem.
theorem 4.1.
Let the assumptions (B.1)(B.3) hold. Then all of the following statements are equivalent:
  1. (i)
    There is a constant $\sigma_{1}>0$ such that
    $\langle h,(\nabla^{2}f(\bar{x})+Q)h\rangle\geq\sigma_{1}\|h\|^{2}\quad\forall~{}h\in\mathrm{aff}(S)=\mathsf{A}.$
  2. (ii)
    There are neighborhoods $U$ of $\bar{x}$ and $V$ of 00 such that for every $v\in V$ there is a point $u\in U\cap(\partial\psi)^{-1}(v)$ such that:
    $\forall~{}x\in U,\quad\psi(x)\geq\psi(u)+\langle v,x-u\rangle+\frac{\sigma_{2}}{2}\|x-u\|^{2}.$
  3. (iii)
    The subdifferential $\partial\psi$ is strongly metrically regular at $\bar{x}$ for 00 and the necessary condition $\mathrm{d}^{2}\psi(\bar{x}|0)(h)\geq 0$ is satisfied for all $h\in\mathbb{R}^{n}$.
  4. (iv)
    The normal map $F^{\tau}_{\mathrm{nor}}$ is strongly metrically regular at $\bar{z}$ for 00 and the second-order necessary condition $\mathrm{d}^{2}\psi(\bar{x}|0)(h)\geq 0$ holds for all $h\in\mathbb{R}^{n}$.
  5. (v)
    The second-order necessary condition $\mathrm{d}^{2}\psi(\bar{x}|0)(h)\geq 0$ holds for all $h\in\mathbb{R}^{n}$ and every matrix $M\in\mathcal{M}_{\mathrm{nor}}^{\tau}(\bar{z})$ is invertible.
  6. (vi)
    The strong second-order necessary condition $\langle h,(\nabla^{2}f(\bar{x})+Q)h\rangle\geq 0$ holds for all $h\in\mathsf{A}$ and $M=\nabla^{2}f(\bar{x})D+\frac{1}{\tau}(I-D)$ is invertible for all $D\in\partial_{B}\mathrm{prox}_{\tau\varphi}(\bar{z})$.
  7. (vii)
    There is a constant $\sigma_{3}>0$ such that:
    $\forall~{}D\in\partial_{B}\mathrm{prox}_{\tau\varphi}(\bar{z}),\quad D\nabla^{2}f(\bar{x})D+\tau^{-1}D(I-D)\succeq\sigma_{3}D^{2}.$
  8. (viii)
    There is a constant $\sigma_{4}>0$ such that:
    $\forall~{}D\in\partial\mathrm{prox}_{\tau\varphi}(\bar{z}),\quad D\nabla^{2}f(\bar{x})D+\tau^{-1}D(I-D)\succeq\sigma_{4}D^{2}.$
  9. (ix)
    There is a neighborhood $U$ of $\bar{z}$ and $\sigma_{5}>0$ such that:
    $\forall~{}z\in U,\,\forall~{}D\in\partial\mathrm{prox}_{\tau\varphi}(z),\quad D\nabla^{2}f(\mathrm{prox}_{\tau\varphi}(z))D+\tau^{-1}D(I-D)\succeq\sigma_{5}D^{2}.$
Each of the conditions (i)(ix) implies the following statement – which is further equivalent to (i)(ix) provided that $I-\tau\nabla^{2}f(\bar{x})$ is invertible:
  • (x)
    The natural residual $F^{\tau}_{\mathrm{nat}}$ is strongly metrically regular at $\bar{x}$ for 00 and the necessary condition $\mathrm{d}^{2}\psi(\bar{x}|0)(h)\geq 0$ is satisfied for all $h\in\mathbb{R}^{n}$.
Condition (i) is the $\mathrm{SSOSC}$ Eq. 8. Condition (ii) is an equivalent characterization of tilt-stability of the local minimizer $\bar{x}$. It is further equivalent to a strong second-order condition based on Mordukhovich’s coderivative, see [14, Theorem 4.6 (ii)]. We refer to [40, 13, 14] for more background and additional relations. Conditions (iii), (iv), and (x) connect the $\mathrm{SSOSC}$ to the strong metric regularity of the subdifferential, the normal map, and the natural residual. Following the discussion after Eq. 9, invertibility of the matrices $M\in\mathcal{M}_{\mathrm{nor}}^{\tau}(\bar{z})$ in condition (v) is equivalent to the CD-regularity of $F^{\tau}_{\mathrm{nor}}$ at $\bar{z}$. Due to $\mathcal{M}_{\mathrm{nat}}^{\tau}(\bar{x})=\tau\mathcal{M}_{\mathrm{nor}}^{\tau}(\bar{z})^{\top}$, condition (v) can also be equivalently expressed using $\mathcal{M}_{\mathrm{nat}}^{\tau}(\bar{x})$ and the natural residual $F^{\tau}_{\mathrm{nat}}$. In particular, noticing $\partial F^{\tau}_{\mathrm{nat}}(\bar{x})h\subseteq\mathcal{M}_{\mathrm{nat}}^{\tau}(\bar{x})h$, $h\in\mathbb{R}^{n}$, (see [8]), condition (v) implies CD-regularity of $F^{\tau}_{\mathrm{nat}}$ at $\bar{x}$. (If $I-\tau\nabla^{2}f(\bar{x})$ is invertible, then it follows $\mathcal{M}_{\mathrm{nat}}^{\tau}(\bar{x})=\partial F^{\tau}_{\mathrm{nat}}(\bar{x})$ and (v) is fully equivalent to CD-regularity of $F^{\tau}_{\mathrm{nat}}$, cf. [7, Lemma 1]). Condition (vi) is a BD-regularity-type analogue of (v). Conditions (vii) and (viii) are the Bouligand and Clarke variant of the second-order condition Eq. 9 that was initially proposed in [36]. The generalized version (ix) of (viii) was used in [36] to prove global-to-local transition and superlinear convergence of a trust region-based semismooth Newton method.
We note that when $\psi$ is convex, then $\mathrm{d}^{2}\psi(\bar{x}|0)$ is convex due to [47, Theorem 13.20], and the strong second-order necessary condition holds automatically. Therefore, in the convex case, we obtain equivalence between BD- and CD-regularity. Similar results have been established for convex semidefinite programming problems [7].
We continue with two counterexamples before starting with the proof of 4.1. First, one might wonder whether the strong second-order necessary condition in (vi) can be relaxed to the regular second-order necessary condition. The following counterexample provides a negative answer.
Example 4.2.
We set $f(x):=\frac{1}{2}\langle x,Ax\rangle$ and $\varphi(x):=\iota_{\mathbb{R}^{2}_{+}}(x)$, where $a_{11}=a_{12}=a_{21}=2$ and $a_{22}=1$. As polyhedral function, $\varphi$ is clearly $C^{2}$-fully decomposable at all $x\in\mathbb{R}_{+}^{2}$, see [49, 31]. Defining $\bar{x}:=(0,0)^{\top}$, we have $\mathrm{d}^{2}\psi(\bar{x}|0)(w)=\langle w,Aw\rangle+\iota_{\mathbb{R}_{+}^{2}}(w)$ for all $w$. It is not hard to see that the second-order necessary condition is satisfied at $\bar{x}$. We may further set $\tau=1$, $\bar{z}=\bar{x}-\tau\nabla f(\bar{x})=\bar{x}$, and $\bar{\lambda}=0$. Elementary calculus shows $\partial_{B}\Pi_{\mathbb{R}_{+}^{2}}(\bar{z})=\{I,B_{1},B_{2},0\}$, where $B_{1}=\mathrm{diag}(1,0)$ and $B_{2}=\mathrm{diag}(0,1)$. The set $\{AD+I-D:D\in\partial_{B}\Pi_{\mathbb{R}_{+}^{2}}(\bar{z})\}$ contains the following four matrices:
$A,\quad\begin{pmatrix}2&0\\ 2&1\end{pmatrix},\quad\begin{pmatrix}1&2\\ 0&1\end{pmatrix},\quad I.$
They are all invertible. But the CD-regularity-type condition (v) fails to hold, since $\mathrm{aff}(\mathbb{R}^{2}_{+})=\mathbb{R}^{2}$ and $A$ is not positive definite.
Next, we present an example to demonstrate that the additional invertibility of $I-\tau\nabla^{2}f(\bar{x})$ is necessary to establish equivalence between (x) and the other conditions.
Example 4.3.
We consider $f(x):=\frac{1}{2}\langle x,Ax\rangle$ and $\varphi(x):=\iota_{K}(x)$, where $A:=\mathrm{diag}(2,-1)$ and $K:=\{x\in\mathbb{R}^{2}:x_{1}\geq|x_{2}|\}$. As in Example 4.2, $\varphi$ is $C^{2}$-fully decomposable at all $x\in K$ and for $\bar{x}=(0,0)^{\top}$, we obtain $\mathrm{d}^{2}\psi(\bar{x}|0)(w)=\langle w,Aw\rangle+\iota_{K}(w)$ for all $w$. As before, the second-order necessary condition $\mathrm{d}^{2}\psi(\bar{x}|0)(h)\geq 0$ is satisfied at $\bar{x}$ for all $h\in\mathbb{R}^{2}$. Moreover, the projection onto $K$ can be computed via:
$\begin{array}[]{llll}\Pi_{K}(x)=\frac{x_{1}+x_{2}}{2}(1,1)^{\top}&\text{if }x_{2}\geq|x_{1}|,&\Pi_{K}(x)=\frac{x_{1}-x_{2}}{2}(1,-1)^{\top}&\text{if }x_{2}\leq-|x_{1}|,\\[4.30554pt] \Pi_{K}(x)=(0,0)^{\top}&\text{if }x_{1}\leq-|x_{2}|,&\Pi_{K}(x)=x&\text{if }x_{1}\geq|x_{2}|.\end{array}$
We now set $\tau=\frac{1}{2}$. Then, we obtain $x-\tau\nabla f(x)=x-\frac{1}{2}Ax=(0,\frac{3}{2}x_{2})^{\top}$ and
$F^{\tau}_{\mathrm{nat}}(x)=x-\Pi_{K}(x-\tau\nabla f(x))=\begin{pmatrix}x_{1}-\frac{3}{4}|x_{2}|&\frac{1}{4}x_{2}\end{pmatrix}^{\top}.$
To show that $F^{\tau}_{\mathrm{nat}}$ is strongly metrically regular at $\bar{x}$ for 00, we calculate its inverse, i.e., we consider the nonlinear equation $F^{\tau}_{\mathrm{nat}}(x)=y$. It readily follows
$(F^{\tau}_{\mathrm{nat}})^{-1}(y)=\begin{pmatrix}y_{1}+3|y_{2}|&4y_{2}\end{pmatrix}^{\top}.$
Therefore, $(F^{\tau}_{\mathrm{nat}})^{-1}$ is single-valued and globally Lipschitz continuous (as it is piecewise linear). Consequently, by [47, Theorem 9.43], $F^{\tau}_{\mathrm{nat}}$ is strongly metrically regular at $\bar{x}$ for 00. However, condition (i) in 4.1 does not hold, since $A$ is not positive semidefinite and $\mathrm{aff}(K)=\mathbb{R}^{2}$.
Next, we state a preparatory result that is used to prove the implication “(viii) $\implies$ (ix)”. The proof of 4.4 applies similar techniques as in [36, Proposition 7.11] and can be found in Appendix B.
Proposition 4.4.
Let $D\in\mathbb{S}^{n}_{+}$, $B\in\mathbb{S}^{n}$ and $\tau,\sigma>0$ be given with:
$DBD+{\tau^{-1}}D(I-D)\succeq\sigma D^{2}.$
Then, there are neighborhoods $U$ of $D$ and $V$ of $B$ such that for every $\tilde{D}\in U\cap\mathbb{S}^{n}_{+}$ and $\tilde{B}\in V\cap\mathbb{S}^{n}$, we have $\tilde{D}\tilde{B}\tilde{D}+{\tau^{-1}}\tilde{D}(I-\tilde{D})\succeq\tfrac{\sigma}{4}\tilde{D}^{2}$.
We now present the proof of 4.1.
Proof of 4.1.
Note that the equivalences (i) $\iff$ (vii) $\iff$ (viii) have been established in Section 2 and Section 3. In fact, we immediately have (i) $\iff$ (viii) and (viii) $\implies$ (vii). To show the remaining implication (vii) $\implies$ (i), we now argue that the matrix $P=\Pi_{\mathsf{A}}(I+\tau Q)^{-1}\Pi_{\mathsf{A}}$ in Eq. P.2 is actually an element of $\partial_{B}\mathrm{prox}_{\tau\varphi}(\bar{z})$ under Eq. P.1. By Eq. P.1 and due to $\mathcal{R}(D)\subseteq\mathsf{A}$, it holds that
$D=\Pi_{\mathcal{R}(D)}(I+\tau A)^{-1}\Pi_{\mathcal{R}(D)}\preceq D=\Pi_{\mathcal{R}(D)}(I+\tau\Pi_{\mathcal{R}(D)}Q\Pi_{\mathcal{R}(D)})^{-1}\Pi_{\mathcal{R}(D)}\preceq P$
for all $D\in\partial\mathrm{prox}_{\tau\varphi}(\bar{z})$. We claim that $P$ is an extreme point of the set $\partial\mathrm{prox}_{\tau\varphi}(\bar{z})=\mathrm{conv}(\partial_{B}\mathrm{prox}_{\tau\varphi}(\bar{z}))$. Let us suppose that this claim is wrong, i.e., there exist $P_{1},P_{2}\in\partial\mathrm{prox}_{\tau\varphi}(\bar{z})$, $P_{1}\neq P_{2}$, and $\mu\in(0,1)$ such that $P=\mu P_{1}+(1-\mu)P_{2}$. Then, it follows
$0\preceq\mu(P-P_{1})=(1-\mu)(P_{2}-P)\preceq 0$
and we can infer $P-P_{1}=P_{2}-P=0$ which yields a contradiction. Hence, $P$ is an extreme point of $\partial\mathrm{prox}_{\tau\varphi}(\bar{z})$ and we have $P\in\partial_{B}\mathrm{prox}_{\tau\varphi}(\bar{z})$. The strategies utilized in the proof of 2.3 are then directly applicable to complete the implication (vii) $\implies$ (i). We continue the proof of 4.1 in a step-by-step fashion:
(ix) $\iff$ (viii): Clearly, we have (ix) $\implies$ (viii). The direction (viii) $\implies$ (ix) follows from 4.4, the local Lipschitz continuity of $\mathrm{prox}_{\tau\varphi}$, and the upper semicontinuity of $\partial\mathrm{prox}_{\tau\varphi}$.
(viii) $\implies$ (v): The second-order necessary condition in (v) follows from (viii) $\iff$ (i). Invertibility of the matrices $M\in\mathcal{M}_{\mathrm{nor}}^{\tau}(\bar{z})$ was verified after introducing Eq. 9.
(v) $\implies$ (iv): As discussed, the invertibility condition in (v) implies CD-regularity of $F^{\tau}_{\mathrm{nor}}$ at $\bar{z}$. Thus, (iv) follows from the classical Clarke inverse function theorem, cf. [8, Theorem 7.1.1] or [12, Theorem 4D.3].
(iv) $\implies$ (iii): By assumption, there exists a neighborhood $U$ of $\bar{z}$ and $\kappa>0$ such that $\|F^{\tau}_{\mathrm{nor}}(z_{1})-F^{\tau}_{\mathrm{nor}}(z_{2})\|\geq\kappa\|z_{1}-z_{2}\|$ for all $z_{1},z_{2}\in U$. In addition, there is a neighborhood $V$ of 00 such that $(F^{\tau}_{\mathrm{nor}})^{-1}:V\to\mathbb{R}^{m}$ is well-defined, single-valued, and Lipschitz continuous, cf. [12, Proposition 3G.1] and we may assume $(F^{\tau}_{\mathrm{nor}})^{-1}(V)\subseteq U$. Let $(x_{1},y_{1}),(x_{2},y_{2})\in(\mathbb{R}^{n}\times V)\cap\mathrm{gph}(\partial\psi)$ be arbitrary. Then, we can write:
$y_{1}=\nabla f(x_{1})+v_{1},\;y_{2}=\nabla f(x_{2})+v_{2}\quad\text{for some}\quad v_{1}\in\partial\varphi(x_{1}),\;v_{2}\in\partial\varphi(x_{2}).$
Thus, setting $z_{1}=x_{1}+\tau v_{1}$, $z_{2}=x_{2}+\tau v_{2}$, we have $x_{1}=\mathrm{prox}_{\tau\varphi}(z_{1})$, $x_{2}=\mathrm{prox}_{\tau\varphi}(z_{2})$ and $y_{1}=F^{\tau}_{\mathrm{nor}}(z_{1})$, $y_{2}=F^{\tau}_{\mathrm{nor}}(z_{2})$. Utilizing $y_{1},y_{2}\in V$, this implies $z_{1}=(F^{\tau}_{\mathrm{nor}})^{-1}(y_{1})$ and $z_{2}=(F^{\tau}_{\mathrm{nor}})^{-1}(y_{2})$ and we can infer $z_{1},z_{2}\in U$. Moreover, this verifies that $(\partial\psi)^{-1}$ is single-valued on $V$, i.e., we have $(\partial\psi)^{-1}(y)=\{\mathrm{prox}_{\tau\varphi}((F^{\tau}_{\mathrm{nor}})^{-1}(y))\}$ for all $y\in V$. In addition, applying the Lipschitz continuity of the proximity operator, it follows
$\kappa_{p}\kappa\|x_{1}-x_{2}\|\leq\kappa\|z_{1}-z_{2}\|\leq\|F^{\tau}_{\mathrm{nor}}(z_{1})-F^{\tau}_{\mathrm{nor}}(z_{2})\|=\|y_{1}-y_{2}\|,$
for some $\kappa_{p}>0$. This establishes the strong metric regularity of $\partial\psi$ at $\bar{x}$ for 00.
(iii) $\implies$ (ii): The necessary condition $\mathrm{d}^{2}\psi(\bar{x}|0)(h)\geq 0$, $h\in\mathbb{R}^{n}$, implies condition (3.6) in [14, Corollary 3.3(i)]; see, e.g., [47, Theorem 13.24]. Therefore, by [14, Corollary 3.3], $\bar{x}$ is a local minimum of $\psi$. The conclusion then follows from [14, Theorem 3.7].
(ii) $\implies$ (i): Following the proof of 3.14 there are sequences $x^{k}\to\bar{x}$ and $\lambda^{k}\to\bar{\lambda}$ such that:
$\mathrm{d}^{2}\varphi(x^{k}|v^{k})(w)=\langle w,H(x^{k},\lambda^{k})w\rangle\quad\forall~{}w\in DF(x^{k})^{-1}\mathrm{aff}(N_{\mathcal{Q}}(\bar{\lambda})),$
where $v^{k}=DF(x^{k})^{\top}\lambda^{k}$. In particular, setting $u^{k}=\nabla f(x^{k})+v^{k}$, this implies
(19) $\mathrm{d}^{2}\psi(x^{k}|u^{k})(w)=\langle w,[\nabla^{2}f(x^{k})+H(x^{k},\lambda^{k})]w\rangle\quad\forall~{}w\in DF(x^{k})^{-1}\mathrm{aff}(N_{\mathcal{Q}}(\bar{\lambda})).$
By [14, Theorem 3.7], $(\partial\psi)^{-1}$ is single-valued around 00. Hence, using $u^{k}\to\nabla f(\bar{x})+\bar{v}=0$ and (ii), it holds that $(\partial\psi)^{-1}(u^{k})=\{x^{k}\}$ and
(20) $\psi(x)\geq\psi(x^{k})+\langle u^{k},x-x^{k}\rangle+\frac{\sigma_{2}}{2}\|x-x^{k}\|^{2}$
for all $x$ in a neighborhood of $\bar{x}$ and all $k$. Let us fix $w\in\mathrm{aff}(S)=DF(\bar{x})^{-1}\mathrm{aff}(N_{\mathcal{Q}}(\bar{\lambda}))$. By 3.11, we can choose $w^{k}\to w$ with $w^{k}\in DF(x^{k})^{-1}\mathrm{aff}(N_{C}(\bar{\lambda}))$ and by the definition of $\mathrm{d}^{2}\psi(x^{k}|u^{k})$, there are $w^{k,i}\to w^{k}$ and $t_{i}\downarrow 0$ such that:
$\mathrm{d}^{2}\psi(x^{k}|u^{k})(w^{k})=\lim_{i\to\infty}\frac{\psi(x^{k}+t_{i}w^{k,i})-\psi(x^{k})-t_{i}\langle u^{k},w^{k,i}\rangle}{\frac{1}{2}t_{i}^{2}}\geq\sigma_{2}\|w^{k}\|^{2},$
where we applied Eq. 20. Thus, using Eq. 19, we can infer $\langle w^{k},[\nabla^{2}f(x^{k})+H(x^{k},\lambda^{k})]w^{k}\rangle\geq\sigma_{2}\|w^{k}\|^{2}$ for all $k$. Taking the limit $k\to\infty$, this shows that the $\mathrm{SSOSC}$ in (i) holds.
(v) $\implies$ (vi) $\implies$ (i): By the previous steps, we have (v) $\iff$ (i) and hence, (v) implies that the strong second-order necessary condition is satisfied. The BD-regularity like condition in (vi) is clearly implied by (v). The direction (vi) $\implies$ (i) is similar to the proof of 2.3 and “(vii) $\implies$ (i)” by using Eq. P.2 and [36, Lemma 6.3].
(v) $\implies$ (x): Noticing the CD-regularity of $F^{\tau}_{\mathrm{nat}}$ at $\bar{x}$, the proof is identical to the direction (v) $\implies$ (iv).
(x) $\implies$ (iv): Suppose that (iv) is not true. Then, we can find sequences $\{z^{k}_{1}\}_{k}$, $\{z^{k}_{2}\}_{k}$ with $z^{k}_{1},z^{k}_{2}\to\bar{z}$, $z_{1}^{k}\neq z_{2}^{k}$, and $\alpha_{k}\to 0$ such that:
(21) $\|F^{\tau}_{\mathrm{nor}}(z^{k}_{1})-F^{\tau}_{\mathrm{nor}}(z^{k}_{2})\|\leq\alpha_{k}\|z^{k}_{1}-z^{k}_{2}\|.$
We set $G(x):=x-\tau\nabla f(x)$. Since $DG(\bar{x})=I-\tau\nabla^{2}f(\bar{x})$ is invertible, $G$ is a local diffeomorphism at $\bar{x}$ satisfying $G(\bar{x})=\bar{z}$. In particular, both $G$ and $G^{-1}$ are well-defined around $\bar{x}$ and $\bar{z}$ and we can define $x_{1}^{k}:=G^{-1}(z_{1}^{k})$ and $x_{2}^{k}:=G^{-1}(z_{2}^{k})$. (Without loss of generality, we assume that $x_{1}^{k}$ and $x_{2}^{k}$ are well-defined in this way for all $k\in\mathbb{N}$). The continuity of $G^{-1}$ further implies $x_{1}^{k},x_{2}^{k}\to\bar{x}$ as $k\to\infty$. Moreover, utilizing the strong metric regularity of $F^{\tau}_{\mathrm{nat}}$ at $\bar{x}$ for 00 and the continuity of $DG$, we may assume
(22) $\|F^{\tau}_{\mathrm{nat}}(x^{k}_{1})-F^{\tau}_{\mathrm{nat}}(x^{k}_{2})\|\geq\kappa\|x^{k}_{1}-x^{k}_{2}\|,\quad\|G(x^{k}_{1})-G(x^{k}_{2})\|\leq L\|x_{1}^{k}-x_{2}^{k}\|$
for all $k$ and for suitable $\kappa,L>0$. In addition, for all $x,z\in\mathbb{R}^{n}$ with $G(x)=z$, we can represent the normal map and natural residual as follows:
$F^{\tau}_{\mathrm{nor}}(z)=\tau^{-1}[G(x)-G(\mathrm{prox}_{\tau\varphi}(z))]\quad\text{and}\quad F^{\tau}_{\mathrm{nat}}(x)=x-\mathrm{prox}_{\tau\varphi}(z).$
Setting $p_{1}^{k}=\mathrm{prox}_{\tau\varphi}(z_{1}^{k})$ and $p^{k}_{2}=\mathrm{prox}_{\tau\varphi}(z^{k}_{2})$, we then obtain
$\tau[F^{\tau}_{\mathrm{nor}}(z_{1}^{k})-F^{\tau}_{\mathrm{nor}}(z_{2}^{k})]$ $=G(x_{1}^{k})-G(x^{k}_{2})-[G(\mathrm{prox}_{\tau\varphi}(z^{k}_{1}))-G(\mathrm{prox}_{\tau\varphi}(z^{k}_{2}))]$
$=B(x_{1}^{k},x^{k}_{2})(x_{1}^{k}-x^{k}_{2})-B(p^{k}_{1},p^{k}_{2})(p^{k}_{1}-p^{k}_{2})$
$=[B(x_{1}^{k},x_{2}^{k})-B(p^{k}_{1},p^{k}_{2}))](x^{k}_{1}-x^{k}_{2})+B(p_{1}^{k},p^{k}_{2})[F^{\tau}_{\mathrm{nat}}(x_{1}^{k})-F^{\tau}_{\mathrm{nat}}(x^{k}_{2})]$
where $B(y,y^{\prime})=\int_{0}^{1}DG(y^{\prime}+t(y-y^{\prime}))\,\mathrm{d}t$. Noticing $\bar{x}=\mathrm{prox}_{\tau\varphi}(\bar{z})$, we further have $B(x_{1}^{k},x_{2}^{k}),B(p_{1}^{k},p_{2}^{k})\to DG(\bar{x})=I-\tau\nabla^{2}f(\bar{x})$ as $k\to\infty$. Hence, there are $\bar{\sigma}>0$ and $K\in\mathbb{N}$ such that $\sigma_{\min}{(B(p^{k}_{1},p^{k}_{2}))}\geq\bar{\sigma}$ and
$\quad\|[B(x_{1}^{k},x_{2}^{k})-B(p^{k}_{1},p^{k}_{2}))](x^{k}_{1}-x^{k}_{2})\|\leq\frac{\kappa\bar{\sigma}}{2}\|x^{k}_{1}-x^{k}_{2}\|$
for all $k\geq K$ which by Eq. 22 implies $\|F^{\tau}_{\mathrm{nor}}(z_{1}^{k})-F^{\tau}_{\mathrm{nor}}(z_{2}^{k})\|\geq\frac{\kappa\bar{\sigma}}{2\tau}\|x_{1}^{k}-x_{2}^{k}\|$. Combining this estimate with Eq. 21 and Eq. 22, we have $\frac{\kappa\bar{\sigma}}{2\tau}\|x_{1}^{k}-x_{2}^{k}\|\leq\|F^{\tau}_{\mathrm{nor}}(z_{1}^{k})-F^{\tau}_{\mathrm{nor}}(z_{2}^{k})\|\leq\alpha_{k}\|z_{1}^{k}-z_{2}^{k}\|\leq L\alpha_{k}\|x_{1}^{k}-x_{2}^{k}\|$ for all $k\geq K$. Due to $\alpha_{k}\to 0$, this yields a contradiction and hence, (iv) is true.
5 Conclusion
We study a strong second-order sufficient condition ($\mathrm{SSOSC}$) for composite-type problems whose nonsmooth part has a generalized conic quadratic second subderivative. The equivalence between the $\mathrm{SSOSC}$ and a second-order condition involving the normal map is established based on algebraic assumptions on the generalized Jacobian of the proximity operator. These structural conditions are then shown to hold for $C^{2}$-strictly decomposable functions under an additional mild geometric assumption on the support set $\mathcal{Q}$. Finally, the equivalence between the $\mathrm{SSOSC}$ and other variational concepts are discussed, including CD-regularity and the strong metric regularity of the subdifferential, the normal map, and the natural residual. The developed theory is applicable to a broad and generic class of conic-type composite problems for which closed-form expressions of the corresponding proximity operator and generalized Jacobians are not necessarily known and can be utilized in the local convergence analysis of the semismooth Newton method.
Appendix A Proof of 1.5
Proof.
To show (i), we note that $G_{\mathrm{ex}}(x)=\bigcap\{\partial\sigma_{C}(z):x\in\partial\sigma_{C}(z)\}$. This automatically implies $N_{C}(x)\subseteq N_{C}(y)$. Let $v\in N_{C}(y)$, by [45, Theorem 18.1], we know $G_{\mathrm{ex}}(x)\subset\partial\sigma_{C}(v)$, which proves $v\in N_{C}(x)$ and $N_{C}(x)=N_{C}(y)$. The equality for the tangent cone follows from $T_{C}(q)=N_{C}(q)^{\circ}$ for all $q\in C$. To show (ii), let us take $z\in\partial\sigma_{C}(y)$. By definition, we have $\langle z-x,y\rangle=\sigma_{C}(y)-\sigma_{C}(y)=0$ and $z-x\in T_{C}(x)=N_{C}(x)^{\circ}$. Consequently, it follows $z-x\in\{y\}^{\perp}\cap T_{C}(x)=N_{N_{C}(x)}(y)=\mathrm{lin}(T_{C}(x))$ by [2, Example 6.40]. Then, for any $v\in N_{C}(x)$, we have $\langle v,x\rangle=\sigma_{C}(v)=\langle v,z\rangle$, which proves that $z\in\partial\sigma_{C}(v)$ and $z\in G_{\mathrm{ex}}(x)$. This yields $G_{\mathrm{ex}}(x)=\partial\sigma_{C}(y)$ since $G_{\mathrm{ex}}(x)\subseteq\partial\sigma_{C}(y)$ follows from the definition.
Appendix B Proof of 4.4
Proof.
Let $\tilde{D}$ be sufficiently close to $D$. According to [53, Lemma 4.3], there exist eigendecompositions of $\tilde{D}$ and $D$ such that:
$\tilde{D}=\tilde{P}^{\top}\tilde{\Lambda}\tilde{P},\quad D=P^{\top}\Lambda P,\quad\|P-\tilde{P}\|=\mathcal{O}(\|D-\tilde{D}\|),\quad\|\Lambda-\tilde{\Lambda}\|=\mathcal{O}(\|D-\tilde{D}\|).$
Here, $\Lambda$ and $\tilde{\Lambda}$ are given by $\Lambda=\operatorname{diag}(\lambda_{1},\dots,\lambda_{n})$ with $\lambda_{1}\geq\dots\geq\lambda_{\ell}>0=\lambda_{\ell+1}=\dots=\lambda_{n}$ and $\tilde{\Lambda}=\operatorname{diag}(\tilde{\lambda}_{1},\dots,\tilde{\lambda}_{n})$ with $\tilde{\lambda}_{1}\geq\dots\geq\lambda_{q}>0=\lambda_{q+1}=\dots=\lambda_{n}$. By the Lipschitz continuity of the eigenvalues [21, Theorem 4.3.1], we clearly have $q\geq\ell$ for all $\tilde{D}$ sufficiently close to $D$. Let us set $A=\operatorname{diag}(\lambda_{1},\dots,\lambda_{\ell})$, $\tilde{A}=\operatorname{diag}(\tilde{\lambda}_{1},\dots,\tilde{\lambda}_{\ell})$, and $A_{q}=\operatorname{diag}(\tilde{\lambda}_{1},\dots,\tilde{\lambda}_{q})$ and let $\Pi_{r}\in\mathbb{R}^{n\times r}$, $\Pi_{r}x:=(x^{\top},0)^{\top}$ be the natural projection onto the first $r$ components. The assumption is equivalent to
$Q:=\Pi_{\ell}^{\top}PBP^{\top}\Pi_{\ell}+\tau^{-1}(A^{-1}-I)\succeq\sigma I.$
Now, let us take sufficiently small neighborhoods $U$ of $D$ and $V$ of $B$, such that for all $\tilde{D}\in U$ and $\tilde{B}\in V$, it holds that $\tilde{Q}:=\Pi_{l}^{\top}\tilde{P}\tilde{B}\tilde{P}^{\top}\Pi_{l}+{\tau^{-1}}(\tilde{A}^{-1}-I)\succeq\tfrac{\sigma}{2}I$.
Let us note the choice of $P$ depends on $\tilde{D}$, but the errors $\|P-\tilde{P}\|$ and $\|A^{-1}-\tilde{A}^{-1}\|$ are uniform for all $\tilde{D}$ sufficiently close to $D$, so the error $\|Q-\tilde{Q}\|$ can be arbitrarily small provided that $\|P-\tilde{P}\|$, $\|A^{-1}-\tilde{A}^{-1}\|$, and $\|B-\tilde{B}\|$ are sufficiently small. The desired condition $\tilde{D}\tilde{B}\tilde{D}+\frac{1}{\tau}\tilde{D}(I-\tilde{D})\succeq\frac{\sigma}{4}\tilde{D}^{2}$ can also be expressed as:
$R:=\Pi_{q}^{\top}\tilde{P}\tilde{B}\tilde{P}^{\top}\Pi_{q}+{\tau^{-1}}(A_{q}^{-1}-I)\succeq\tfrac{\sigma}{4}I.$
Next, let us introduce $L_{1}:=\mathrm{span}\{e_{1},\dots,e_{\ell}\}$ and $L_{2}:=\mathrm{span}\{e_{\ell+1},\dots,e_{q}\}$. For every $h\in\mathbb{R}^{q}$, we can write $h=h_{1}+h_{2}$ with $h_{1}\in L_{1}$ and $h_{2}\in L_{2}$. It then follows:
$h^{\top}Rh$ $=h_{1}^{\top}Rh_{1}+2h_{1}^{\top}Rh_{2}+h_{2}^{\top}Rh_{2}$
$=h_{1}^{\top}\tilde{Q}h_{1}+2h_{1}^{\top}Rh_{2}+h_{2}^{\top}Rh_{2}\geq\tfrac{\sigma}{2}\|h_{1}\|^{2}-2|h_{1}^{\top}Rh_{2}|+h_{2}^{\top}Rh_{2}.$
If $q=\ell$, then we are done since $h=h_{1}$. Now assume $q>\ell$. Then, we have:
$h_{2}^{\top}Rh_{2}=h_{2}^{\top}\Pi_{q}^{\top}\tilde{P}\tilde{B}\tilde{P}^{\top}\Pi_{q}h_{2}+{\tau^{-1}}h_{2}^{\top}(A_{q}^{-1}-I)h_{2}\geq[{\tau^{-1}}(\tilde{\lambda}_{\ell+1}^{-1}-1)-\|\tilde{B}\|]\|h_{2}\|^{2},$
$|h_{1}^{\top}Rh_{2}|=|h_{1}^{\top}\Pi_{q}^{\top}\tilde{P}\tilde{B}\tilde{P}^{\top}\Pi_{q}h_{2}|\leq\|\tilde{B}\|\|h_{1}\|\|h_{2}\|$, and by Young’s inequality, we can infer $2\|\tilde{B}\|\|h_{1}\|\|h_{2}\|\leq\frac{\sigma}{4}\|h_{1}\|^{2}+\frac{4\|\tilde{B}\|^{2}}{\sigma}\|h_{2}\|^{2}$. Therefore, we obtain
$h^{\top}Rh\geq\tfrac{\sigma}{4}\|h_{1}\|^{2}+\left[{\tau^{-1}}(\tilde{\lambda}_{\ell+1}^{-1}-1)-\|\tilde{B}\|-\tfrac{4\|\tilde{B}\|^{2}}{\sigma}\right]\|h_{2}\|^{2}.$
Shrinking $U,V$ (if necessary), we can ensure ${\tau^{-1}}(\tilde{\lambda}_{\ell+1}^{-1}-1)-\|\tilde{B}\|-\frac{4\|\tilde{B}\|^{2}}{\sigma}\geq\frac{\sigma}{4}$, which finishes the proof.